### v1

# Knowledge Sharing and Transfer via Centralized Reward Agent for Multi-Task Reinforcement Learning

Abstract
Reward shaping is effective in addressing the sparse-reward challenge in reinforcement learning by providing immediate feedback through auxiliary informative rewards. Based on the reward shaping strategy, we propose a novel multi-task reinforcement learning framework, that integrates a centralized reward agent (CRA) and multiple distributed policy agents. The CRA functions as a knowledge pool, which aims to distill knowledge from various tasks and distribute it to individual policy agents to improve learning efficiency. Specifically, the shaped rewards serve as a straightforward metric to encode knowledge. This framework not only enhances knowledge sharing across established tasks but also adapts to new tasks by transferring valuable reward signals. We validate the proposed method on both discrete and continuous domains, demonstrating its robustness in multi-task sparse-reward settings and its effective transferability to unseen tasks.
1 Introduction
Reinforcement learning (RL) has made significant progress across various domains, like robotics Kober et al., (2013), gaming Lample and Chaplot, (2017), autonomous vehicles Aradi, (2020), signal processing Luo et al., (2024), and large language models Shinn et al., (2023). However, environments with sparse and delayed rewards remain a significant challenge, as the absence of immediate feedback hinders the agent from distinguishing the value of states and leads to aimless exploration (Ladosz et al.,, 2022). Reward Shaping (RS) has been proven to be an effective technique to address this challenge by providing additional dense and informative rewards Sorg et al., 2010b ; Sorg et al., 2010a . Concurrently, multi-task reinforcement learning (MTRL) is becoming increasingly important for its ability to share and transfer knowledge across tasks. In this context, the auxiliary rewards infused with task-specific information in RS offer a straightforward entry to distribute knowledge among different tasks. Therefore, we believe that integrating RS techniques into MTRL is a highly promising and intuitive direction to enhance the efficacy of multi-task system learning.
Presently, numerous MTRL algorithms involving knowledge transfer have been developed. Policy distillation methods identify and combine the commonalities across different policies Rusu et al., (2016); Teh et al., (2017); Parisotto et al., (2016); Xu et al., (2024); representation sharing methods extract and disseminate the common features or gradients among agents Yang et al., (2020); D’Eramo et al., (2020); Sodhani et al., (2021); and parameter sharing methods design architectural modules to reuse parameters or layers across networks Sun et al., (2022); Cheng et al., (2023). Despite their potential, these strategies often encounter challenges such as slow adaptation and response to the transferred knowledge, or suffer from delays in reutilizing and comprehending this information. Therefore, leveraging the RS mechanism, which directly adds a metric to the reward function, offers a promising alternative to address these limitations.
Regarding the RS techniques, not all shaped rewards serve effectively as mediums for knowledge transfer. Specifically, the intrinsic motivation based rewards are typically designed using heuristics to generate task-agnostic signals. Examples include incorporating exploration bonuses Bellemare et al., (2016); Ostrovski et al., (2017); Devidze et al., (2022), rewarding novel states Tang et al., (2017); Burda et al., (2018), and encouraging curiosity-driven behaviors Pathak et al., (2017); Mavor-Parker et al., (2022). Although these approaches broaden extensive exploration, they are not directly related to specific tasks and thus lack transferability. Consequently, we focus on another branch of RS methods, the task-contextual rewards, which automatically learn and encode task-specific information, such as the hidden values, contributions of states, or future-oriented insights, that can be shared effectively across various tasks Ma et al., 2024b ; Mguni et al., (2023); Memarian et al., (2021); Ma et al., 2024a .
To share task-related knowledge in MTRL via RS techniques, and inspired by the ReLara framework Ma et al., 2024b , which integrates an assistant reward agent to densify sparse environmental rewards, we propose the Centralized Reward Agent based MTRL framework (CenRA) [^1] The source code is accessible at: https://github.com/mahaozhe/CenRA. The framework consists of two main components: a centralized reward agent (CRA) and multiple distributed policy agents. Each policy agent individually learns control behaviors within their respective tasks and shares their experiences with the CRA, which are collected from interactions with their environment. The CRA extracts common knowledge from these experiences and learns to generate dense rewards that are encoded with task-specific information. These rewards are then distributed back to the policy agents to augment their original environmental rewards. Additionally, given that different tasks may contribute variably to the MTRL system, we introduce an information synchronization mechanism to further balance the knowledge distribution, considering task similarity and agent learning progress, to ensure a system-wide optimal performance. The main contributions of this paper are summarized as follows:
  1. (i)
    We propose the CenRA framework targeted to solve MTRL problems. It incorporates a CRA that functions as a centralized knowledge pool, efficiently distilling and distributing valuable information from various tasks to policy agents, and further adapting to new tasks.
  2. (ii)
    CenRA leverages the advantage of reward shaping techniques to infuse insights via dense rewards. This approach not only provides a direct metric for policy agents to absorb knowledge but also effectively addresses the sparse-reward challenge.
  3. (iii)
    We introduce an information synchronization mechanism that considers both task similarity and agent convergence progress to balance multi-task learning. This mechanism provides a novel direction for maintaining system equilibrium in MTRL.
  4. (iv)
    CenRA is validated in both discrete and continuous control MTRL environments, where the extrinsic rewards are sparse. CenRA outperforms several baseline models in terms of learning efficiency, knowledge transferability, and system-wide performance.
2 Related Work
Multi-task reinforcement learning (MTRL) has attracted significant attention recently due to its potential to share knowledge across multiple tasks, thereby improving learning performance Caruana, (1993). We discuss existing MTRL literature from three main directions:
Knowledge Transfer methods focus on identifying and transferring task-relevant features across diverse tasks Zeng et al., (2021). Policy distillation Rusu et al., (2016) is a well-studied approach to extract and share task-specific behaviors or representations that many works are built on: Teh et al., (2017) introduced Distral, which distills a centroid policy from multiple task-policies; Parisotto et al., (2016) developed Actor-Mimic, where a single policy is trained to mimic several expert policies from different tasks; while Yin and Pan, (2017) incorporated hierarchical prioritized experience replay buffer to select and learn multi-task experiences; Hessel et al., (2019) further proposed an adaptation mechanism to equalize the impact of each task in policy distillation. Additionally, Xu et al., (2020) explored the transfer of offline knowledge to train policies, and further leveraged online learning for fine-tuning. Bai et al., (2023) introduced a dual-phase learning approach, optimizing individual policies while correcting them across multiple tasks. Mysore et al., (2022) used separate critics for each task to accompany a single actor to integrate their feedback. These methods mitigate gradient interference to an extent, however, balancing the distribution of knowledge across tasks is crucial. Without a careful trade-off, the performance of the entire system could be compromised.
Representation Sharing methods explore architectural solutions of reusing network modules or representing commonalities to the MTRL problem D’Eramo et al., (2020); Devin et al., (2017); Ma et al., 2024c ; Hong et al., (2021). Sun et al., (2022) used a parameter compositional approach to learn and share a subspace of parameters, allowing policies for various tasks to be interpolated within it. Yang et al., (2020) employed soft modularization to learn foundational policies and utilized a routing network to generate probabilities to combine them. He et al., (2024) introduced the Dynamic Depth Routing framework which dynamically adjusts the use of network modules in response to task difficulty. Sodhani et al., (2021) leveraged task-related metadata to create composable representations. Cheng et al., (2023) and Lan et al., (2023) both incorporated attention mechanisms: the former employed attention-based mixture of experts to capture task relationships, while the latter used Temporal Attention for contrastive learning purposes. Although these methods demonstrate efficacy in learning shared representations, they may struggle to fully capture the complexity of highly diverse tasks. Moreover, adapting shared structures to new tasks typically requires extra design efforts.
Single-Policy Generalization methods learn a single policy to solve multiple tasks simultaneously or continuously, in the absence of information from prior policies or task-specific details, in which case, the primary goal is to enhance the policy’s generalization capabilities. Model-free meta-learning techniques have been proposed to enhance the multi-task generalization Finn et al., (2017). Yang et al., (2017) designed a sharing network structure that allows an agent to learn multiple tasks concurrently. Vuong et al., (2019) introduced a confidence-sharing agent to detect and define shared regions between tasks to support single policy learning. Wan et al., (2020) proposed a transfer learning framework to handle mismatches in state and action spaces. Additionally, several methods focus on overcoming gradient interference to enhance the generalization in various tasks Chen et al., (2018); Yu et al., (2020), while Ammar et al., (2014) developed a consecutive learning policy gradient approach. These methods are efficient in saving computational resources, but the generalization ability of the policy may be constrained when faced with out-of-distribution or previously unseen tasks.
3 Preliminaries
Markov Decision Process (MDP) fundamentally models sequential decision-making problems under uncertainty. An MDP represents the interaction between an agent and its environment as a tuple $\langle S,A,P,R,\gamma\rangle$, where $S$ is the state space, $A$ is the action space, $P:S\times A\times S\rightarrow[0,1]$ is the probability of transiting from one state to another given an action, $R:S\times A\rightarrow\mathbb{R}$ is the reward function, and $\gamma\in[0,1]$ is the discount factor to modulate the importance of future versus immediate reward.
Multi-Task Reinforcement Learning (MTRL) addresses the challenge of learning multiple tasks simultaneously within an integrated model to leverage commonalities and differences across tasks. Typically, MTRL introduces a task space $\mathcal{T}$, assuming all tasks are sampled from this space and thus follow a unique distribution. Each task is modeled as an independent MDP. An MTRL agent aims to learn optimal policies $\pi_{i}:S\rightarrow A$ for each task $T_{i}\sim\mathcal{T}$, to maximize their corresponding expected cumulative rewards, or returns, denoted by $G_{i}=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}R_{i}(s_{t},a_{t})]$.
RL with an Assistant Reward Agent (ReLara) Ma et al., 2024b introduces a dual-agent framework designed to tackle the challenge of sparse rewards in RL. The operational architecture of ReLara is shown in Figure 1. Within this framework, the original agent is termed as policy agent, while an assistant reward agent is integrated to enrich the feedback mechanism by generating dense, informative rewards. The reward agent, trained as a self-contained RL agent, autonomously extracts hidden value information from the environmental states (and potentially the actions of the policy agent) to craft meaningful reward signals. These signals significantly improve learning efficiency by providing immediate and pertinent bonuses.
Figure 1: The ReLara framework Ma et al., 2024b .
4 Methodology
We extend the ReLara framework Ma et al., 2024b , which utilizes a specialized reward agent to extract environmental information and densify the sparse rewards, to MTRL scenarios. Our proposed framework is termed as Centralized Reward Agent based MTRL (CenRL), which incorporates a centralized reward agent (CRA) to support simultaneous reinforcement learning agents across multiple tasks. A high-level illustration of the CenRA framework is shown in Figure 2. The CRA is responsible for extracting general task-specific knowledge from various tasks and distributing valuable information to the policy agents by reconstructing their reward models. The detailed methodology for knowledge extraction and sharing is introduced in Section 4.1. Furthermore, we recognize there are potential disparities in the information that each task could contribute, which might lead to an imbalance in knowledge distribution. To mitigate this, we introduce an information synchronization mechanism by considering two main factors: the similarity of the tasks and the online learning performance of the policy agents, details given in Section 4.2. Finally, the overall framework of CenRA is presented in Section 4.3.
Figure 2: A high-level illustration of the CenRA framework.
4.1 Knowledge Extraction and Share
4.1.1 Problem Formulation
We consider an MTRL setting comprising $N$ distinct tasks $\{T_{1},T_{2},\dots,T_{N}\}$, all executed within the same type of environment $\mathcal{E}$. We assume that the dimensionality and structure of state space $S$ and action space $A$ remain uniform across tasks, to ensure the centralized reward agent processes consistent inputs. Despite this uniformity, each task may feature unique state sets, goals, and transition dynamics. For instance, a series of mazes with the same size and shape but varying map configurations and objects would satisfy this condition. For the environment $\mathcal{E}$, we denote states by $s\in S$ and actions by $a\in A$. Additionally, for each task $T_{i}$, we denote the transition function as $P_{i}(s^{\prime}|s,a)$ and the reward function as $R_{i}(s,a)$.
The CenRA framework consists of a centralized reward agent $\mathcal{A}^{rwd}$ and multiple policy agents $\{\mathcal{A}^{pol}_{1},\allowbreak\mathcal{A}^{pol}_{2},\dots,\mathcal{A}^{pol}_{N}\}$. Each policy agent $\mathcal{A}^{pol}_{i}$ operates independently to complete its corresponding task $T_{i}$, utilizing appropriate RL algorithms as backbones based on the specific characteristics of the task. For example, implementing DQN Mnih et al., (2015) for discrete control decisions, while TD3 Fujimoto et al., (2018) or SAC Haarnoja et al., 2018a for continuous control tasks. Moreover, the policy of the reward agent $\mathcal{A}^{rwd}$ is denoted as $\pi^{rwd}$, and the internal policy of policy agent $\mathcal{A}^{pol}_{i}$ is denoted as $\pi^{pol}_{i}$.
4.1.2 Centralized Reward Agent
The CRA $\mathcal{A}^{rwd}$ aims to extract environment-relevant knowledge and distribute it to policy agents by generating additional dense rewards to support their original reward functions. Similar to the ReLara framework Ma et al., 2024b , we model the CRA as a self-contained RL agent, yet, as an extension to ReLara, our CRA is designed to concurrently interact with multiple policy agents and their respective tasks. The CRA’s policy $\pi^{rwd}$ generates continuous rewards given both an environmental state and a policy agent’s behavior. Specifically, $\pi^{rwd}$ maps the Cartesian product of the state space and action space, $S\times A$, to a defined reward space, which constrains the rewards to a range of real numbers, $\mathcal{R}=[R_{min},R_{max}]\subset\mathbb{R}$. For simplicity’s sake, we denote the observation of CRA as $s^{rwd}=(s_{i},a_{i})$, where $s_{i}\sim T_{i}$ and $a_{i}\sim\pi^{pol}_{i}(s_{i})$. To distinguish from the environmental reward, the generated reward is termed as knowledge reward, denoted as $r^{knw}$.
We adopt an off-policy actor-critic algorithm to optimize the CRA Konda and Tsitsiklis, (1999). To aggregate and reuse experiences from all policy agents, a concatenated replay buffer $\mathcal{D}=\bigcup_{i=1}^{N}{\mathcal{D}_{i}}$ is constructed, where $\mathcal{D}_{i}$ represents the replay buffer of each policy agent $\mathcal{A}^{pol}_{i}$. Besides, each transition is augmented with the CRA-generated knowledge reward, $r^{knw}$. Specifically, the transition from policy agent $\mathcal{A}^{pol}_{i}$ stored in the replay buffer is defined as $\tau=(s^{rwd}_{t},r^{knw}_{t},r^{env}_{t},s^{rwd}_{t+1}|T_{i})$. The augmented transition includes all necessary information for optimizing both the CRA and each corresponding policy agent, thus making the concatenated replay buffer a shared resource across the entire framework and minimizing storage overhead.
The CRA’s update process involves using these stored transitions to optimize the reward-generating actor $\pi^{rwd}$ and the value estimation critic. The temporal difference error is given by:
$\delta_{t}=r^{env}_{t}+\gamma V^{rwd}(s^{rwd}_{t+1})-V^{rwd}(s^{rwd}_{t})|T_{i},$
where $\tau_{t}=(s^{rwd}_{t},r^{env}_{t},s^{rwd}_{t+1}|T_{i})\sim\mathcal{D}$. The objective function for optimizing the critic module is formulated as:
$J(V^{rwd})=\underset{\tau_{t}\sim\mathcal{D}}{\mathbb{E}}[\delta_{t}^{2}],$
which aims to minimize the squared TD error. Concurrently, the actor module is updated through the following objective function:
$J(\pi^{rwd})\!=\!\underset{\tau_{t}\sim\mathcal{D}}{\mathbb{E}}\Big{[}\underset{r^{knw}_{t}\sim\pi^{rwd}(\cdot|s^{rwd}_{t})}{\mathbb{E}}\big{[}\log\pi^{rwd}(r^{knw}_{t}|s^{rwd}_{t})\cdot\delta_{t}\big{]}\Big{]}.$
4.1.3 Policy Agents with Knowledge Rewards
Each policy agent $\mathcal{A}^{pol}_{i}$ interacts with its specific environment and stores the experiences in its corresponding replay buffer $\mathcal{D}_{i}$. Policy agents receive two types of rewards: the environmental reward $r^{env}_{i}$ from their respective task $T_{i}$ and the knowledge reward $r^{knw}$ from CRA. The augmented reward is given by:
$r^{pol}_{i}=r^{env}_{i}+\lambda r^{knw},\quad r^{knw}\sim\pi^{rwd}(\cdot|s_{i},a_{i}),$
where $\lambda\in(0,1]$ is a scaling weight factor. The optimal policy ${\pi^{pol}_{i}}^{*}$ for each agent is derived by maximizing the cumulative augmented reward:
${\pi^{pol}_{i}}^{*}=\operatorname*{arg\,max}_{\pi^{pol}_{i}}\underset{(s_{i},a_{i})\sim\pi^{pol}_{i}}{\mathbb{E}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}r^{pol}_{i}\Big{]}.$
It is worth noting that the environmental reward $r^{env}_{i}$ is retrieved from the replay buffer (if adopting an off-policy approach). Conversely, the knowledge reward $r^{knw}$ is computed in real-time using the most recently updated $\mathcal{A}^{rwd}$, ensuring it reflects the latest learning advancements. Lastly, each policy agent is able to employ any suitable RL algorithm, whether on-policy or off-policy, to best address its specific task, which enhances the CenRA framework’s generality and flexibility.
4.2 Information Synchronization of Policy Agents
In the CenRA framework, the information provided by different tasks may exhibit significant disparities, potentially leading to an imbalance in knowledge extraction and distribution. In this section, we introduce an information synchronization mechanism for CenRA to maintain a balanced manner from the perspective of the entire system. Specifically, this mechanism is implemented by controlling the quantity of samples that CRA retrieves from each task’s replay buffer $\mathcal{D}_{i}$ by a sampling weight $\bm{w}$. We mainly consider two aspects: the similarity among tasks and the real-time learning performance of the policy agents.
Similarity Weight is derived from the similarity among tasks, enabling the CRA to focus on relatively outlier tasks. To simplify computation, we use the hidden layers extracted from each policy agent’s neural network encoders to represent the tasks’ features. To reduce randomness, we average the hidden features of the most recent $K$ steps. We adopt a cross-attention mechanism, which is widely used in neural networks, to calculate the similarity weight Vaswani et al., (2017). Specifically, for task $T_{i}$, let $\bm{H}_{i}$ denote the averaged hidden feature vector, which serves as the key, and the centroid of all tasks $\bm{c}$ acts as the query. Then, the similarity $s_{i}$ of task $T_{i}$ to the centroid of the task cluster is calculated as:
$s_{i}=\frac{\bm{c}^{T}\cdot\bm{H}_{i}}{\sqrt{D}},\quad\bm{c}=\frac{1}{N}\sum_{k=1}^{N}\bm{H}_{k},$
where $D$ is the dimension of the hidden feature to prevent gradient vanishing or exploding. A larger $s_{i}$ indicates a greater similarity between $T_{i}$ and the centroid. Given our assumption is that the tasks farther from the centroid require more attention, the similarity weight is defined as $\bm{w}^{sim}=\text{Softmax}\big{(}[1/s_{1},1/s_{2},\dots,1/s_{N}]\big{)}$.
Performance Weight is determined by the real-time learning performance of each policy agent, to ensure the CRA focuses more on lagging tasks. Similar to the similarity weight, we average the environmental rewards $r^{env}_{i}$ from the most recent $K$ steps, denoted as $R^{tail}_{i}$, to measure the recent learning trends. The performance weight is then defined as $\bm{w}^{per}=\text{Softmax}\big{(}[1/R^{tail}_{1},1/R^{tail}_{2},\dots,1/R^{tail}_{N}]\big{)}$.
Lastly, the final sampling weight $\bm{w}$ combines the similarity weight and the performance weight as $\bm{w}=\alpha\bm{w}^{sim}+(1-\alpha)\bm{w}^{per}$, where $\alpha$ is a hyperparameter to balance the two aspects. The CRA samples from each replay buffer $\mathcal{D}_{i}$ according to $\bm{w}$, ensuring a balanced and effective knowledge extraction and learning.
4.3 Overall Framework
The overall framework of CenRA is summarized in Algorithm 1. The CRA and policy agents are updated alternately and asynchronously, with the frequency of updating the CRA adjustable according to the actual situation. Sampling weights are calculated in real-time, using the most recently optimized encoders and the current learning performance, ensuring CRA continuously adjusts its focus to optimally balance knowledge extraction across multiple tasks.
The learned CRA acts as a robust knowledge pool, which is able to support new tasks as an additional module by transferring knowledge through auxiliary reward signals. This is particularly beneficial in sparse-reward environments, as the knowledge rewards can guide the policy agents toward the correct direction and reduce the exploration burden. Additionally, the CRA can be further optimized alongside new tasks in a continuous learning scheme that enhances adaptability and effectiveness in dynamic settings.
Algorithm 1 Centralized Reward Agent based MTRL
0:  Multiple tasks $\{T_{1},T_{2},\dots,T_{N}\}$.
0:  Policy agents $\{\mathcal{A}^{pol}_{1},\mathcal{A}^{pol}_{2},\dots,\mathcal{A}^{pol}_{N}\}$.
0:  Centralized reward agent $\mathcal{A}^{rwd}$.
0:  Concatenated replay buffer $\mathcal{D}=\bigcup_{i=1}^{N}{\mathcal{D}_{i}}$.
1:  for each iteration do
2:     for each task $T_{i}$ do
3:        $(s_{t},a_{t},r^{env}_{t},s_{t+1},a_{t+1})\sim\text{Interact}(\mathcal{A}^{pol}_{i},T_{i})$triangleright Interact and collect one transition
4:        $r^{knw}_{t}\sim\mathcal{A}^{rwd}(s_{t},a_{t})$triangleright Sample an off-policy knowledge reward
5:        $s^{rwd}_{t}=(s_{t},a_{t})$, $s^{rwd}_{t+1}=(s_{t+1},a_{t+1})$
6:        $\mathcal{D}_{i}\leftarrow\mathcal{D}_{i}\cup\{(s^{rwd}_{t},r^{knw}_{t},r^{env}_{t},s^{rwd}_{t+1}|T_{i})\}$triangleright Store the transition in corresponding $\mathcal{D}_{i}$
7:        Update policy agent $\mathcal{A}^{pol}_{i}$triangleright Update $\mathcal{A}^{pol}_{i}$ using backbone RL algorithm
8:     end for
9:     Get similarity and performance weights $\bm{w}^{sim}$, $\bm{w}^{per}$
10:     $\bm{w}=\alpha\bm{w}^{sim}+(1-\alpha)\bm{w}^{per}$
11:     $\{s^{rwd}_{t},r^{knw}_{t},r^{env}_{t},s^{rwd}_{t+1}|T_{i}\}_{\mathcal{B}}\sim\mathcal{D}|\bm{w}$triangleright Draw a batch of samples based on the sampling weight
12:     Update centralized reward agent $\mathcal{A}^{rwd}$
13:  end for
5 Experiments
We conduct experiments in three environments with multiple tasks: 2DMaze and 3DPickup by Chevalier-Boisvert et al., (2024), and MujocoCar by Ji et al., (2023), which include both discrete and continuous control problems. All tasks provide challenging sparse environmental rewards, where the agent only receives a reward of $1$ upon successful completion of the final objective, and a reward of 00 is given in all other states. Detailed configurations are shown in Figure 3, with each task’s setting provided in Appendix A.
(a) 2DMaze
(b) 3DPickup
(c) MujocoCar
Figure 3: Experimental environments with multiple tasks. (a) 2DMaze: A $10\times 10$ 2D maze environment where the agent must pick up a key and then pass through a door to exit. Different tasks involve different placements of walls, keys, and doors. (b) 3DPickup: A 3D environment where the agent’s task is to navigate to and pick up target objects. Tasks vary by different object types, e.g., balls, cubes, keys and health kits, and their locations. (c) MujocoCar: A mujoco-based race car aims to navigate to specified areas, where each task has a different target location.
5.1 Comparative Evaluation in MTRL
In this section, we evaluate the learning performance of CenRA in the MTRL context, configuring each environment with four different tasks. We benchmark CenRA against several state-of-the-art baselines, including: (a) the backbone algorithms of the policy agents in CenRA: the Deep Q-Network (DQN) Mnih et al., (2015) for the discrete control tasks in 2DMaze and 3DPickup, and the Soft Actor-Critic (SAC) Haarnoja et al., 2018b for the continuous control tasks in MujocoCar; (b) The RL with an Assistant Reward Agent (ReLara) algorithm Ma et al., 2024b , which learns an additional reward agent to generate informative rewards. ReLara can be considered as a special case of CenRA where each policy agent is paired with a separate reward agent, without cross-task information sharing; (c) The Policy Optimization and Policy Correction (PiCor) algorithm Bai et al., (2023), which uses policy correction to adapt and transfer information across policies among individual tasks; (d) The Multi-Critic Actor Learning (MCAL) algorithm Mysore et al., (2022), which maintains separate critics for each task for value estimation and knowledge representation, while training a unified multi-task actor. We use the CleanRL library Huang et al., (2022) or the author-provided codes to implement these baselines.
(a) 2DMaze
(b) 3DPickup
(c) MujocoCar
Figure 4: Comparison of the learning performance of CenRA with the baselines.
Table 1: Comparison of episodic returns and standard errors.
Algo. 2DMaze
Maze #1 Maze #2 Maze #3 Maze #4
CenRA 0.89 $\pm$ 0.03 0.91 $\pm$ 0.02 0.92 $\pm$ 0.02 0.93 $\pm$ 0.02
DQN 0.77 $\pm$ 0.05 0.49 $\pm$ 0.09 0.54 $\pm$ 0.09 0.78 $\pm$ 0.05
ReLara 0.83 $\pm$ 0.05 0.88 $\pm$ 0.03 0.76 $\pm$ 0.05 0.74 $\pm$ 0.05
PiCor 0.80 $\pm$ 0.05 0.85 $\pm$ 0.05 0.78 $\pm$ 0.06 0.84 $\pm$ 0.05
MCAL 0.76 $\pm$ 0.07 0.72 $\pm$ 0.09 0.57 $\pm$ 0.08 0.69 $\pm$ 0.08
Algo. 3DPickup
Ball Cube Key Health kit
CenRA 0.95 $\pm$ 0.02 0.68 $\pm$ 0.09 0.80 $\pm$ 0.06 0.69 $\pm$ 0.07
DQN 0.34 $\pm$ 0.06 0.07 $\pm$ 0.03 0.28 $\pm$ 0.03 0.28 $\pm$ 0.07
ReLara 0.45 $\pm$ 0.08 0.39 $\pm$ 0.10 0.33 $\pm$ 0.09 0.29 $\pm$ 0.08
PiCor 0.28 $\pm$ 0.07 0.44 $\pm$ 0.08 0.35 $\pm$ 0.09 0.28 $\pm$ 0.10
MCAL 0.24 $\pm$ 0.07 0.34 $\pm$ 0.05 0.36 $\pm$ 0.08 0.45 $\pm$ 0.07
Algo. MujocoCar
Target #1 Target #2 Target #3 Target #4
CenRA 588.2 $\pm$ 0.73 549.3 $\pm$ 0.64 447.7 $\pm$ 0.67 474.3 $\pm$ 0.66
SAC 247.3 $\pm$ 0.53 210.6 $\pm$ 0.49 177.7 $\pm$ 0.41 156.4 $\pm$ 0.38
ReLara 434.3 $\pm$ 0.68 436.0 $\pm$ 0.67 308.1 $\pm$ 0.58 420.8 $\pm$ 0.69
PiCor 411.0 $\pm$ 0.66 426.0 $\pm$ 0.63 470.7 $\pm$ 0.66 442.5 $\pm$ 0.70
MCAL 341.0 $\pm$ 0.62 315.6 $\pm$ 0.58 430.4 $\pm$ 0.59 389.8 $\pm$ 0.59
(a) Maze #1
(b) Maze #2
(c) Maze #3
(d) Maze #4
Figure 5: Visualization of the action yielding the maximum knowledge rewards in the four 2DMaze tasks.
We train multiple instances with different random seeds for each task to report the average data. The results of the episodic returns and their standard errors throughout training are shown in Figure 4 and Table 1. To ensure a fair comparison, we use the same hyperparameters (if applicable) and network structures across all experiments, the details are provided in Appendix B.
We observe that CenRA consistently outperforms the baselines in three main aspects. First, CenRA achieves the highest episodic returns in nearly all tasks, demonstrating superior learning efficiency and faster convergence. Importantly, it also maintains good stability and robustness, which can be seen in the significantly fewer fluctuations and oscillations especially after convergence, compared to other models. Second, unlike PiCor and MCAL, which have shown uneven progress across different tasks in the same environment, CenRA maintains a well-balanced performance by showing relatively consistent learning progress and minimal variability in learning outcomes across each four-task group. This ensures that no single task’s performance is disproportionately high at the expense of others, which is crucial in multi-task learning. Third, the centralized reward agent in CenRA effectively enhances knowledge sharing among tasks. This is evident from the comparison with ReLara, which uses independent reward agents and lacks the mechanism for knowledge exchange. By extracting and distributing insights from one task to another, the CRA improves the learning efficiency of individual tasks, highlighting the advantages of integrated knowledge management.
5.2 What Has the Centralized Reward Agent Learned?
In this section, we visualize the learned knowledge rewards by the centralized reward agent $\mathcal{A}^{rwd}$ in the 2DMaze environment. After training on the four tasks in Section 5.1 of the paper, we let the CRA generate the knowledge rewards for each action in every state and visualize the action direction that yields the maximum rewards, $a^{*}=\operatorname*{arg\,max}_{a}{{\pi^{rwd}}^{*}(s_{i},a)},s_{i}\sim S$, in Figure 5.
The shaded areas in the figures represent regions within the real task that the agent cannot reach, as it cannot access the space behind the door without picking up the key. However, we forced the agent into these areas for evaluation. Outside the shaded regions, we observe that the CRA successfully learned meaningful knowledge rewards. Before picking up the key, the agent received the highest reward in the corresponding state when moving towards the key. Similarly, after picking up the key, the agent received the highest reward when moving towards the door and the final target. This demonstrates that in scenarios where the original environmental rewards are sparse, these detailed knowledge rewards can effectively guide the agent to converge more quickly.
5.3 Knowledge Transfer to New Tasks
In this section, we assess the centralized reward agent’s ability to transfer learned knowledge to new tasks. We continue using the CRA model trained on the four tasks in Section 5.1 but initiate a new policy agent to tackle a newly designed task that the models have never encountered before. For the CenRA framework, we explore two scenarios: one where the CRA continues to optimize while cooperating with the new policy agent (CenRA w/ learning), and another where only the policy agent is trained without further updates to the CRA (CenRA w/o learning). We compare the two approaches with the backbone algorithms (DQN and SAC) and ReLara. In the ReLara model, the reward agent starts training anew (without pre-learned knowledge). The results are shown in Figure 6 and Table 2.
We observe that CenRA with further learning achieved rapid convergence, mainly due to the CRA’s retention of previously acquired knowledge and its ability to quickly adapt to new tasks through continuous optimization. Notably, even without any additional training, CenRA outperforms both the ReLara, which requires training a new reward agent, and the backbone algorithms, which lack additional information. In the CenRA framework, the CRA stores environment-relevant knowledge and transfers it to other policy agents, thus improving their learning efficiency. Additionally, sharing pre-learned information with policy agents is particularly crucial in our experiments with challenging sparse rewards. Without any external knowledge, learning would require extensive exploration. However, the CRA provides meaningful dense rewards that significantly accelerate the learning process, even during the initial phases.
Figure 6: Comparison of the learning performance of CenRA with the baselines in new tasks.
Table 2: Comparison of the episodic returns and standard errors in new tasks.
Algo. New 2DMaze New 3DPickup New MujocoCar
CenRA w/ learning 0.95 $\pm$ 0.01 0.96 $\pm$ 0.00 532.00 $\pm$ 0.61
CenRA w/o learning 0.89 $\pm$ 0.03 0.68 $\pm$ 0.00 524.73 $\pm$ 0.60
ReLara 0.76 $\pm$ 0.06 0.26 $\pm$ 0.00 224.65 $\pm$ 0.49
DQN/SAC 0.26 $\pm$ 0.08 0.16 $\pm$ 0.00 129.06 $\pm$ 0.30
(a) Before getting the key.
(b) After getting the key.
Figure 7: Actions yielding the highest knowledge rewards at each state in a new 2DMaze task.
To further demonstrate CenRA’s transferability, we select the 2DMaze environment to visualize the learned knowledge of CRA in an unseen task. As shown in Figure 7, we plot the directions of actions that yield the maximum knowledge reward at each position, categorized into two scenarios before and after obtaining the key. While some guidance in peripheral areas may be slightly off-target, most states receive reasonable rewards that align with human understanding. This demonstrates the effective transfer of knowledge and with such dense rewards, the agent’s adaptation to new tasks is able to be well supported.
5.4 Effect of Sampling Weight
We conducted experiments to understand the effects of the information synchronization mechanism in the CenRA framework. Specifically, we compare the full CenRA model against three variants: (a) CenRA without the similarity weight $\bm{w}^{sim}$, (b) CenRA without the performance weight $\bm{w}^{per}$, and (c) CenRA without the entire sampling weight. The comparison results are shown in Figure 8 and Table 3.
(a) 2DMaze
(b) 3DPickup
(c) MujocoCar
Figure 8: Comparison of CenRA with ablation of different batch sampling control weights.
Table 3: Comparison of CenRA with ablation of different batch sampling control weights.
Algo. 2DMaze Var. $\downarrow$ (×$\times 10^{-2}$)
Maze #1 Maze #2 Maze #3 Maze #4
CenRA 0.89 $\pm$ 0.03 0.91 $\pm$ 0.02 0.92 $\pm$ 0.02 0.93 $\pm$ 0.02 0.02
w/o $\bm{w}^{sim}$ 0.88 $\pm$ 0.03 0.92 $\pm$ 0.02 0.87 $\pm$ 0.04 0.82 $\pm$ 0.04 0.17
w/o $\bm{w}^{per}$ 0.76 $\pm$ 0.06 0.88 $\pm$ 0.03 0.82 $\pm$ 0.05 0.87 $\pm$ 0.02 0.28
w/o both 0.63 $\pm$ 0.05 0.83 $\pm$ 0.04 0.63 $\pm$ 0.08 0.80 $\pm$ 0.05 1.24
Algo. 3DPickup Var. $\downarrow$ (×$\times 10^{-2}$)
Ball Cube Key Health kit
CenRA 0.95 $\pm$ 0.02 0.68 $\pm$ 0.09 0.80 $\pm$ 0.06 0.69 $\pm$ 0.07 1.57
w/o $\bm{w}^{sim}$ 0.82 $\pm$ 0.07 0.70 $\pm$ 0.09 0.70 $\pm$ 0.07 0.89 $\pm$ 0.04 0.89
w/o $\bm{w}^{per}$ 0.78 $\pm$ 0.07 0.40 $\pm$ 0.09 0.63 $\pm$ 0.08 0.44 $\pm$ 0.10 3.05
w/o both 0.81 $\pm$ 0.07 0.46 $\pm$ 0.06 0.48 $\pm$ 0.08 0.37 $\pm$ 0.08 3.80
Algo. MujocoCar Var. $\downarrow$ (×$\times 10^{3}$)
Target #1 Target #2 Target #3 Target #4
CenRA 588.2 $\pm$ 0.73 549.3 $\pm$ 0.64 447.7 $\pm$ 0.67 474.3 $\pm$ 0.66 4.24
w/o $\bm{w}^{sim}$ 319.9 $\pm$ 0.59 486.8 $\pm$ 0.71 332.5 $\pm$ 0.70 260.9 $\pm$ 0.54 9.29
w/o $\bm{w}^{per}$ 320.9 $\pm$ 0.89 355.3 $\pm$ 0.68 344.1 $\pm$ 0.87 308.2 $\pm$ 0.72 0.46
w/o both 57.5$\pm$0.35 257.0 $\pm$ 0.68 677.3 $\pm$ 0.54 77.6 $\pm$ 0.26 82.7
The results indicate that the two weights, which control the allocation of samples drawn from each policy agent’s experiences, mainly influence the overall learning performance. Specifically, the absence of sampling weight leads to unbalanced learning outcomes, which is observed by the increased variance in episodic returns across four tasks. While the full CenRA model does not always achieve the lowest variance, it consistently outperforms the other three ablation models regarding overall system performance.
Both weights are essential for the information synchronization mechanism, with the performance weight $\bm{w}^{per}$ having a more important effect. This weight enables CRA to pay more attention to the policy agents that are underperforming or progressing slowly. This suggests that considering the overall learning performance of the multi-task system is also an important objective CenRA seeks to achieve.
6 Discussion and Conclusion
We propose a novel framework named CenRA that innovatively integrates reward shaping into MTRL. The framework not only can share environment-relevant knowledge across tasks to improve learning efficiency, but also effectively addresses the sparse-reward challenge. Specifically, a CRA functions as a knowledge pool, responsible for knowledge distillation and distribution across multiple tasks. Furthermore, our information synchronization mechanism mitigates knowledge distribution disparities across tasks, ensuring optimal system-wide performance. Experiments demonstrate that the dense knowledge rewards generated by the CRA efficiently guide the policy agents’ learning, leading to faster convergence compared to baseline methods. The CenRA framework proves to be robust and performs superiorly in both multi-task learning and new tasks adaptation.
CenRA’s current limitation is its requirement for uniform state and action spaces across tasks. Future improvements could explore preprocessing techniques to adapt the framework for varying state and action spaces to broaden CenRA’s applicability. Additionally, the fixed trade-off between similarity weight and performance weight may not be ideal. A more flexible approach, potentially through adaptive regulation of these weights, could enhance the framework further. Besides, the performance weight might favor underperforming tasks, achieving overall balance but potentially limiting the upper bound of top-performing tasks. This issue could necessitate a better mechanism for trade-off.
7 Acknowledgement
Appendix A Mutli-Task Experimental Configurations
In Section 5 Experiments, we conduct experiments in three environments with multiple tasks: 2DMaze, 3DPickup, and MujocoCar. The detailed configurations of each task are illustrated in Figure 9.
Figure 9: The configurations of multiple tasks used in your experiments.
Appendix B Network Structures and Hyperparameters
B.1 Network Structures
Figure 10 illustrates the structures of all networks employed in our experiments.
Figure 10: The structures of neural networks in our implementation.
B.2 Hyperparameters
We have observed that CenRA demonstrated high robustness and was not sensitive to hyperparameter choices. Table 4 shows the hyperparameters we used in all the experiments.
Table 4: The hyperparameters of CenRA used in our experiments .
Module Hyperparameters Values
Centralized Reward Agent $\mathcal{A}^{rwd}$ discounted factor $\gamma$ 0.99
batch size 256
actor module learning rate $3\times 10^{-4}$
critic module learning rate $1\times 10^{-3}$
policy networks update frequency (steps) $2$
target networks update frequency (steps) $1$
target networks soft update weight $\tau$ $5\times 10^{-3}$
burn-in steps $5000$
Policy Agent $\mathcal{A}^{pol}_{i}$ (DQN Agent) knowledge reward weight $\lambda$ 0.5
discounted factor $\gamma$ 0.99
replay buffer size $|\mathcal{D}_{i}|$ $1\times 10^{6}$
batch size 128
burn-in steps $10000$
Policy Agent $\mathcal{A}^{pol}_{i}$ (SAC Agent) knowledge reward weight $\lambda$ 0.5
discounted factor $\gamma$ 0.99
replay buffer size $|\mathcal{D}_{i}|$ $1\times 10^{6}$
batch size 256
actor module learning rate $3\times 10^{-4}$
critic module learning rate $1\times 10^{-3}$
SAC entropy term factor $\alpha$ learning rate $1\times 10^{-4}$
policy networks update frequency (steps) $2$
target networks update frequency (steps) $1$
target networks soft update weight $\tau$ $5\times 10^{-3}$
burn-in steps $10000$

### v2

@@
-Knowledge Sharing and Transfer via Centralized Reward Agent for Multi-Task Reinforcement Learning
+Centralized Reward Agent for Knowledge Sharing and Transfer in Multi-Task Reinforcement Learning
 Abstract
 Reward shaping is effective in addressing the sparse-reward challenge in reinforcement learning by providing immediate feedback through auxiliary informative rewards.
-Based on the reward shaping strategy, we propose a novel multi-task reinforcement learning framework, that integrates a centralized reward agent (CRA) and multiple distributed policy agents.
+Based on the reward shaping strategy, we propose a novel multi-task reinforcement learning framework that integrates a centralized reward agent (CRA) and multiple distributed policy agents.
 The CRA functions as a knowledge pool, which aims to distill knowledge from various tasks and distribute it to individual policy agents to improve learning efficiency.
 Specifically, the shaped rewards serve as a straightforward metric to encode knowledge.
-This framework not only enhances knowledge sharing across established tasks but also adapts to new tasks by transferring valuable reward signals.
-We validate the proposed method on both discrete and continuous domains, demonstrating its robustness in multi-task sparse-reward settings and its effective transferability to unseen tasks.
+This framework not only enhances knowledge sharing across established tasks but also adapts to new tasks by transferring meaningful reward signals.
+We validate the proposed method on both discrete and continuous domains, including the representative meta world benchmark, demonstrating its robustness in multi-task sparse-reward settings and its effective transferability to unseen tasks.
 1 Introduction
-Reinforcement learning (RL) has made significant progress across various domains, like robotics Kober et al., (2013), gaming Lample and Chaplot, (2017), autonomous vehicles Aradi, (2020), signal processing Luo et al., (2024), and large language models Shinn et al., (2023).
+Reinforcement learning (RL) has made significant progress across various domains, like robotics (Kober et al.,, 2013), gaming (Lample and Chaplot,, 2017), autonomous vehicles (Aradi,, 2020), signal processing (Luo et al.,, 2024), and large language models (Shinn et al.,, 2023; Ouyang et al.,, 2022).
 However, environments with sparse and delayed rewards remain a significant challenge, as the absence of immediate feedback hinders the agent from distinguishing the value of states and leads to aimless exploration (Ladosz et al.,, 2022).
-Reward Shaping (RS) has been proven to be an effective technique to address this challenge by providing additional dense and informative rewards Sorg et al., 2010b ; Sorg et al., 2010a .
+Reward Shaping (RS) has been proven to be an effective technique to address this challenge by providing additional dense and informative rewards (Sorg et al., 2010b, ; Sorg et al., 2010a, ).
 Concurrently, multi-task reinforcement learning (MTRL) is becoming increasingly important for its ability to share and transfer knowledge across tasks.
@@
 Presently, numerous MTRL algorithms involving knowledge transfer have been developed.
-Policy distillation methods identify and combine the commonalities across different policies Rusu et al., (2016); Teh et al., (2017); Parisotto et al., (2016); Xu et al., (2024); representation sharing methods extract and disseminate the common features or gradients among agents Yang et al., (2020); D’Eramo et al., (2020); Sodhani et al., (2021); and parameter sharing methods design architectural modules to reuse parameters or layers across networks Sun et al., (2022); Cheng et al., (2023).
+Policy distillation methods identify and combine the commonalities across different policies (Rusu et al.,, 2016; Teh et al.,, 2017; Parisotto et al.,, 2016; Xu et al.,, 2024; Ma et al., 2024b, ; Ma et al.,, 2023); representation sharing methods extract and disseminate the common features or gradients among agents (Yang et al.,, 2020; D’Eramo et al.,, 2020; Sodhani et al.,, 2021); and parameter sharing methods design architectural modules to reuse parameters or layers across networks (Sun et al.,, 2022; Cheng et al.,, 2023).
 Despite their potential, these strategies often encounter challenges such as slow adaptation and response to the transferred knowledge, or suffer from delays in reutilizing and comprehending this information.
-Therefore, leveraging the RS mechanism, which directly adds a metric to the reward function, offers a promising alternative to address these limitations.
-Regarding the RS techniques, not all shaped rewards serve effectively as mediums for knowledge transfer.
+Therefore, leveraging the RS mechanism, which directly adds a metric to the reward function, offers a compelling alternative to address these limitations.
+Regarding the RS techniques, not all shaped rewards serve effectively as the medium for knowledge transfer.
 Specifically, the intrinsic motivation based rewards are typically designed using heuristics to generate task-agnostic signals.
-Examples include incorporating exploration bonuses Bellemare et al., (2016); Ostrovski et al., (2017); Devidze et al., (2022), rewarding novel states Tang et al., (2017); Burda et al., (2018), and encouraging curiosity-driven behaviors Pathak et al., (2017); Mavor-Parker et al., (2022).
+Examples include incorporating exploration bonuses (Bellemare et al.,, 2016; Ostrovski et al.,, 2017; Devidze et al.,, 2022), rewarding novel states (Tang et al.,, 2017; Burda et al.,, 2018), and encouraging curiosity-driven behaviors (Pathak et al.,, 2017; Mavor-Parker et al.,, 2022).
 Although these approaches broaden extensive exploration, they are not directly related to specific tasks and thus lack transferability.
-Consequently, we focus on another branch of RS methods, the task-contextual rewards, which automatically learn and encode task-specific information, such as the hidden values, contributions of states, or future-oriented insights, that can be shared effectively across various tasks Ma et al., 2024b ; Mguni et al., (2023); Memarian et al., (2021); Ma et al., 2024a .
-To share task-related knowledge in MTRL via RS techniques, and inspired by the ReLara framework Ma et al., 2024b , which integrates an assistant reward agent to densify sparse environmental rewards, we propose the Centralized Reward Agent based MTRL framework (CenRA) [^1] The source code is accessible at: https://github.com/mahaozhe/CenRA.
+Consequently, we focus on another branch of RS methods, the task-contextual rewards, which automatically learn and encode task-specific information, such as the hidden values, contributions of states, or future-oriented insights, that can be shared effectively across various tasks (Ma et al., 2024a, ; Ma et al., 2025b, ; Ma et al., 2025a, ; Mguni et al.,, 2023; Memarian et al.,, 2021).
+To share task-related knowledge in MTRL via RS techniques, and inspired by the ReLara framework (Ma et al., 2024a, ), which integrates an assistant reward agent to densify sparse environmental rewards, we propose the Centralized Reward Agent based MTRL fRAmework (CenRA).
 The framework consists of two main components: a centralized reward agent (CRA) and multiple distributed policy agents.
-Each policy agent individually learns control behaviors within their respective tasks and shares their experiences with the CRA, which are collected from interactions with their environment.
+Each policy agent individually learns control behaviors within their respective tasks and shares their experiences with the CRA.
 The CRA extracts common knowledge from these experiences and learns to generate dense rewards that are encoded with task-specific information.
@@
 2 Related Work
-Multi-task reinforcement learning (MTRL) has attracted significant attention recently due to its potential to share knowledge across multiple tasks, thereby improving learning performance Caruana, (1993).
+Multi-task reinforcement learning (MTRL) has attracted significant attention recently due to its potential to share knowledge across multiple tasks, thereby improving learning performance (Caruana,, 1993).
 We discuss existing MTRL literature from three main directions:
-Knowledge Transfer methods focus on identifying and transferring task-relevant features across diverse tasks Zeng et al., (2021).
-Policy distillation Rusu et al., (2016) is a well-studied approach to extract and share task-specific behaviors or representations that many works are built on: Teh et al., (2017) introduced Distral, which distills a centroid policy from multiple task-policies; Parisotto et al., (2016) developed Actor-Mimic, where a single policy is trained to mimic several expert policies from different tasks; while Yin and Pan, (2017) incorporated hierarchical prioritized experience replay buffer to select and learn multi-task experiences; Hessel et al., (2019) further proposed an adaptation mechanism to equalize the impact of each task in policy distillation.
+Knowledge Transfer methods focus on identifying and transferring task-relevant features across diverse tasks (Zeng et al.,, 2021).
+Policy distillation (Rusu et al.,, 2016) is a well-studied approach to extract and share task-specific behaviors or representations that many works are built on: Teh et al., (2017) introduced Distral, which distills a centroid policy from multiple task-policies; Parisotto et al., (2016) developed Actor-Mimic, where a single policy is trained to mimic several expert policies from different tasks; while Yin and Pan, (2017) incorporated hierarchical prioritized experience replay buffer to select and learn multi-task experiences; Hessel et al., (2019) further proposed an adaptation mechanism to equalize the impact of each task in policy distillation.
 Additionally, Xu et al., (2020) explored the transfer of offline knowledge to train policies, and further leveraged online learning for fine-tuning.
@@
 Without a careful trade-off, the performance of the entire system could be compromised.
-Representation Sharing methods explore architectural solutions of reusing network modules or representing commonalities to the MTRL problem D’Eramo et al., (2020); Devin et al., (2017); Ma et al., 2024c ; Hong et al., (2021).
+Representation Sharing methods explore architectural solutions of reusing network modules or representing commonalities to the MTRL problem (D’Eramo et al.,, 2020; Devin et al.,, 2017; Hong et al.,, 2021).
 Sun et al., (2022) used a parameter compositional approach to learn and share a subspace of parameters, allowing policies for various tasks to be interpolated within it.
 Yang et al., (2020) employed soft modularization to learn foundational policies and utilized a routing network to generate probabilities to combine them.
-He et al., (2024) introduced the Dynamic Depth Routing framework which dynamically adjusts the use of network modules in response to task difficulty.
+He et al., (2024) introduced the Dynamic Depth Routing framework, which dynamically adjusts the use of network modules in response to task difficulty.
 Sodhani et al., (2021) leveraged task-related metadata to create composable representations.
@@
 Single-Policy Generalization methods learn a single policy to solve multiple tasks simultaneously or continuously, in the absence of information from prior policies or task-specific details, in which case, the primary goal is to enhance the policy’s generalization capabilities.
-Model-free meta-learning techniques have been proposed to enhance the multi-task generalization Finn et al., (2017).
+Model-free meta-learning techniques have been proposed to enhance the multi-task generalization (Finn et al.,, 2017).
 Yang et al., (2017) designed a sharing network structure that allows an agent to learn multiple tasks concurrently.
@@
 Wan et al., (2020) proposed a transfer learning framework to handle mismatches in state and action spaces.
-Additionally, several methods focus on overcoming gradient interference to enhance the generalization in various tasks Chen et al., (2018); Yu et al., (2020), while Ammar et al., (2014) developed a consecutive learning policy gradient approach.
+Additionally, several methods focus on overcoming gradient interference to enhance the generalization in various tasks (Chen et al.,, 2018; Yu et al., 2020a, ), while Ammar et al., (2014) developed a consecutive learning policy gradient approach.
 These methods are efficient in saving computational resources, but the generalization ability of the policy may be constrained when faced with out-of-distribution or previously unseen tasks.
 3 Preliminaries
-Markov Decision Process (MDP) fundamentally models sequential decision-making problems under uncertainty.
+Markov Decision Process (MDP) models sequential decision-making problems under uncertainty.
 An MDP represents the interaction between an agent and its environment as a tuple $\langle S,A,P,R,\gamma\rangle$, where $S$ is the state space, $A$ is the action space, $P:S\times A\times S\rightarrow[0,1]$ is the probability of transiting from one state to another given an action, $R:S\times A\rightarrow\mathbb{R}$ is the reward function, and $\gamma\in[0,1]$ is the discount factor to modulate the importance of future versus immediate reward.
@@
 An MTRL agent aims to learn optimal policies $\pi_{i}:S\rightarrow A$ for each task $T_{i}\sim\mathcal{T}$, to maximize their corresponding expected cumulative rewards, or returns, denoted by $G_{i}=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}R_{i}(s_{t},a_{t})]$.
-RL with an Assistant Reward Agent (ReLara) Ma et al., 2024b introduces a dual-agent framework designed to tackle the challenge of sparse rewards in RL.
+RL with an Assistant Reward Agent (ReLara) (Ma et al., 2024a, ) introduces a dual-agent framework designed to tackle the challenge of sparse rewards in RL.
 The operational architecture of ReLara is shown in Figure 1.
@@
 These signals significantly improve learning efficiency by providing immediate and pertinent bonuses.
-Figure 1: The ReLara framework Ma et al., 2024b .
+Figure 1: The ReLara framework Ma et al., 2024a .
 4 Methodology
-We extend the ReLara framework Ma et al., 2024b , which utilizes a specialized reward agent to extract environmental information and densify the sparse rewards, to MTRL scenarios.
-Our proposed framework is termed as Centralized Reward Agent based MTRL (CenRL), which incorporates a centralized reward agent (CRA) to support simultaneous reinforcement learning agents across multiple tasks.
+We propose Centralized Reward Agent fRAmework (CenRA) for MTRL, which incorporates a centralized reward agent (CRA) to support simultaneous reinforcement learning agents across multiple tasks.
 A high-level illustration of the CenRA framework is shown in Figure 2.
@@
 The detailed methodology for knowledge extraction and sharing is introduced in Section 4.1.
-Furthermore, we recognize there are potential disparities in the information that each task could contribute, which might lead to an imbalance in knowledge distribution.
-To mitigate this, we introduce an information synchronization mechanism by considering two main factors: the similarity of the tasks and the online learning performance of the policy agents, details given in Section 4.2.
+Furthermore, to mitigate the potential disparities in the information that each task could contribute, which might lead to an imbalance in knowledge distribution, we introduce an information synchronization mechanism by considering two main factors: the similarity of the tasks and the online learning performance of the policy agents, details given in Section 4.2.
 Finally, the overall framework of CenRA is presented in Section 4.3.
@@
 We consider an MTRL setting comprising $N$ distinct tasks $\{T_{1},T_{2},\dots,T_{N}\}$, all executed within the same type of environment $\mathcal{E}$.
-We assume that the dimensionality and structure of state space $S$ and action space $A$ remain uniform across tasks, to ensure the centralized reward agent processes consistent inputs.
-Despite this uniformity, each task may feature unique state sets, goals, and transition dynamics.
-For instance, a series of mazes with the same size and shape but varying map configurations and objects would satisfy this condition.
-For the environment $\mathcal{E}$, we denote states by $s\in S$ and actions by $a\in A$.
-Additionally, for each task $T_{i}$, we denote the transition function as $P_{i}(s^{\prime}|s,a)$ and the reward function as $R_{i}(s,a)$.
-The CenRA framework consists of a centralized reward agent $\mathcal{A}^{rwd}$ and multiple policy agents $\{\mathcal{A}^{pol}_{1},\allowbreak\mathcal{A}^{pol}_{2},\dots,\mathcal{A}^{pol}_{N}\}$.
-Each policy agent $\mathcal{A}^{pol}_{i}$ operates independently to complete its corresponding task $T_{i}$, utilizing appropriate RL algorithms as backbones based on the specific characteristics of the task.
-For example, implementing DQN Mnih et al., (2015) for discrete control decisions, while TD3 Fujimoto et al., (2018) or SAC Haarnoja et al., 2018a for continuous control tasks.
-Moreover, the policy of the reward agent $\mathcal{A}^{rwd}$ is denoted as $\pi^{rwd}$, and the internal policy of policy agent $\mathcal{A}^{pol}_{i}$ is denoted as $\pi^{pol}_{i}$.
+We assume that the shape of state $s\in S$ and action $a\in A$ remain uniform across tasks, to ensure the CRA processes consistent inputs.
+Despite this uniformity, each task may feature different state spaces, action spaces, goals, and transition dynamics.
+For instance, a series of mazes with the same size but varying map configurations would satisfy this condition.
+For each task $T_{i}$, we denote the transition function as $P_{i}(s^{\prime}|s,a)$ and the reward function as $R_{i}(s,a)$.
+The centralized reward agent (CRA) is denoted as $\mathcal{A}^{rwd}$ and multiple policy agents are denoted as $\{\mathcal{A}^{pol}_{1},\allowbreak\mathcal{A}^{pol}_{2},\dots,\mathcal{A}^{pol}_{N}\}$.
+Each policy agent $\mathcal{A}^{pol}_{i}$ operates independently to complete its corresponding task $T_{i}$, utilizing appropriate RL algorithms as backbones.
+For example, implementing DQN (Mnih et al.,, 2015) for discrete control tasks, while TD3 (Fujimoto et al.,, 2018) or SAC (Haarnoja et al., 2018a, ) for continuous control tasks.
+Moreover, the policy of CRA $\mathcal{A}^{rwd}$ is $\pi^{rwd}$, and the internal policy of policy agent $\mathcal{A}^{pol}_{i}$ is $\pi^{pol}_{i}$, respectively.
 4.1.2 Centralized Reward Agent
 The CRA $\mathcal{A}^{rwd}$ aims to extract environment-relevant knowledge and distribute it to policy agents by generating additional dense rewards to support their original reward functions.
-Similar to the ReLara framework Ma et al., 2024b , we model the CRA as a self-contained RL agent, yet, as an extension to ReLara, our CRA is designed to concurrently interact with multiple policy agents and their respective tasks.
+Similar to the ReLara framework Ma et al., 2024a , we model the CRA as a self-contained RL agent, yet, as an extension to ReLara, our CRA is designed to concurrently interact with multiple policy agents and their respective tasks.
 The CRA’s policy $\pi^{rwd}$ generates continuous rewards given both an environmental state and a policy agent’s behavior.
@@
 The augmented transition includes all necessary information for optimizing both the CRA and each corresponding policy agent, thus making the concatenated replay buffer a shared resource across the entire framework and minimizing storage overhead.
-The CRA’s update process involves using these stored transitions to optimize the reward-generating actor $\pi^{rwd}$ and the value estimation critic. The temporal difference error is given by:
-$\delta_{t}=r^{env}_{t}+\gamma V^{rwd}(s^{rwd}_{t+1})-V^{rwd}(s^{rwd}_{t})|T_{i},$
-where $\tau_{t}=(s^{rwd}_{t},r^{env}_{t},s^{rwd}_{t+1}|T_{i})\sim\mathcal{D}$. The objective function for optimizing the critic module is formulated as:
-$J(V^{rwd})=\underset{\tau_{t}\sim\mathcal{D}}{\mathbb{E}}[\delta_{t}^{2}],$
-which aims to minimize the squared TD error. Concurrently, the actor module is updated through the following objective function:
-$J(\pi^{rwd})\!=\!\underset{\tau_{t}\sim\mathcal{D}}{\mathbb{E}}\Big{[}\underset{r^{knw}_{t}\sim\pi^{rwd}(\cdot|s^{rwd}_{t})}{\mathbb{E}}\big{[}\log\pi^{rwd}(r^{knw}_{t}|s^{rwd}_{t})\cdot\delta_{t}\big{]}\Big{]}.$
+The CRA’s update process involves using these stored transitions to optimize the reward-generating actor $\pi^{rwd}$ and the value estimation critic. The objective function for the critic module is:
+$J(V^{rwd})=\underset{\tau_{t}\sim\mathcal{D}}{\mathbb{E}}[\delta_{t}^{2}],\quad\delta_{t}=r^{env}_{t}+\gamma V^{rwd}(s^{rwd}_{t+1})-V^{rwd}(s^{rwd}_{t})|T_{i},$ (1)
+where $\tau_{t}=(s^{rwd}_{t},r^{env}_{t},s^{rwd}_{t+1}|T_{i})\sim\mathcal{D}$. Concurrently, the actor module is updated through the following objective function:
+$J(\pi^{rwd})=\underset{\tau_{t}\sim\mathcal{D}}{\mathbb{E}}\Big{[}\underset{r^{knw}_{t}\sim\pi^{rwd}(\cdot|s^{rwd}_{t})}{\mathbb{E}}\big{[}\log\pi^{rwd}(r^{knw}_{t}|s^{rwd}_{t})\cdot\delta_{t}\big{]}\Big{]}.$ (2)
 4.1.3 Policy Agents with Knowledge Rewards
@@
 The augmented reward is given by:
-$r^{pol}_{i}=r^{env}_{i}+\lambda r^{knw},\quad r^{knw}\sim\pi^{rwd}(\cdot|s_{i},a_{i}),$
+$r^{pol}_{i}=r^{env}_{i}+\lambda r^{knw},\quad r^{knw}\sim\pi^{rwd}(\cdot|s_{i},a_{i}),$ (3)
 where $\lambda\in(0,1]$ is a scaling weight factor. The optimal policy ${\pi^{pol}_{i}}^{*}$ for each agent is derived by maximizing the cumulative augmented reward:
-${\pi^{pol}_{i}}^{*}=\operatorname*{arg\,max}_{\pi^{pol}_{i}}\underset{(s_{i},a_{i})\sim\pi^{pol}_{i}}{\mathbb{E}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}r^{pol}_{i}\Big{]}.$
+${\pi^{pol}_{i}}^{*}=\operatorname*{arg\,max}_{\pi^{pol}_{i}}\underset{(s_{i},a_{i})\sim\pi^{pol}_{i}}{\mathbb{E}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}r^{pol}_{i}\Big{]}.$ (4)
 It is worth noting that the environmental reward $r^{env}_{i}$ is retrieved from the replay buffer (if adopting an off-policy approach).
@@
 Then, the similarity $s_{i}$ of task $T_{i}$ to the centroid of the task cluster is calculated as:
-$s_{i}=\frac{\bm{c}^{T}\cdot\bm{H}_{i}}{\sqrt{D}},\quad\bm{c}=\frac{1}{N}\sum_{k=1}^{N}\bm{H}_{k},$
+$s_{i}=\frac{\bm{c}^{T}\cdot\bm{H}_{i}}{\sqrt{D}},\quad\bm{c}=\frac{1}{N}\sum_{k=1}^{N}\bm{H}_{k},$ (5)
 where $D$ is the dimension of the hidden feature to prevent gradient vanishing or exploding.
@@
 8:     end for
-9:     Get similarity and performance weights $\bm{w}^{sim}$, $\bm{w}^{per}$
-10:     $\bm{w}=\alpha\bm{w}^{sim}+(1-\alpha)\bm{w}^{per}$
-11:     $\{s^{rwd}_{t},r^{knw}_{t},r^{env}_{t},s^{rwd}_{t+1}|T_{i}\}_{\mathcal{B}}\sim\mathcal{D}|\bm{w}$triangleright Draw a batch of samples based on the sampling weight
-12:     Update centralized reward agent $\mathcal{A}^{rwd}$
-13:  end for
+9:     $\bm{w}=\alpha\bm{w}^{sim}+(1-\alpha)\bm{w}^{per}$triangleright Calculate sampling weight
+10:     $\{s^{rwd}_{t},r^{knw}_{t},r^{env}_{t},s^{rwd}_{t+1}|T_{i}\}_{\mathcal{B}}\sim\mathcal{D}|\bm{w}$triangleright Draw samples based on the sampling weight
+11:     Update centralized reward agent $\mathcal{A}^{rwd}$
+12:  end for
 5 Experiments
-We conduct experiments in three environments with multiple tasks: 2DMaze and 3DPickup by Chevalier-Boisvert et al., (2024), and MujocoCar by Ji et al., (2023), which include both discrete and continuous control problems.
-All tasks provide challenging sparse environmental rewards, where the agent only receives a reward of $1$ upon successful completion of the final objective, and a reward of 00 is given in all other states.
-Detailed configurations are shown in Figure 3, with each task’s setting provided in Appendix A.
-(a) 2DMaze
-(b) 3DPickup
-(c) MujocoCar
-Figure 3: Experimental environments with multiple tasks.
-(a) 2DMaze: A $10\times 10$ 2D maze environment where the agent must pick up a key and then pass through a door to exit.
-Different tasks involve different placements of walls, keys, and doors.
-(b) 3DPickup: A 3D environment where the agent’s task is to navigate to and pick up target objects.
-Tasks vary by different object types, e.g., balls, cubes, keys and health kits, and their locations.
-(c) MujocoCar: A mujoco-based race car aims to navigate to specified areas, where each task has a different target location.
+We conduct experiments in four MTRL domains as shown in Figure 3: the widely used Meta-World benchmark (including ML10 with 10 tasks and ML50 with 50 tasks) (Yu et al., 2020b, ), 2DMaze, 3DPickup (Chevalier-Boisvert et al.,, 2024), and MujocoCar (Ji et al.,, 2023).
+All tasks, including those in Meta-World, are crafted to provide sparse environmental rewards, where the agent receives a reward of $1$ only upon successful completion of the final objective, and 00 otherwise.
+The detailed task configurations are provided in Appendix A.
+(a) Meta-World (ML10-sparse and ML50-sparse)
+(b) 2DMaze
+(c) 3DPickup
+(d) MujocoCar
+Figure 3: Environments with multiple tasks.
+(a) Meta-World: two sparse-reward versions are used: ML10-sparse and ML50-sparse, including diverse robotic manipulation tasks.
+(b) 2DMaze: 2D maze tasks where the agent must pick up a key and then pass through a door to exit.
+(b) 3DPickup: 3D maze tasks where the agent aims to navigate to and pick up different target objects at different locations.
+(c) MujocoCar: mujoco-based race car aims to navigate to different specified areas.
 5.1 Comparative Evaluation in MTRL
-In this section, we evaluate the learning performance of CenRA in the MTRL context, configuring each environment with four different tasks.
-We benchmark CenRA against several state-of-the-art baselines, including: (a) the backbone algorithms of the policy agents in CenRA: the Deep Q-Network (DQN) Mnih et al., (2015) for the discrete control tasks in 2DMaze and 3DPickup, and the Soft Actor-Critic (SAC) Haarnoja et al., 2018b for the continuous control tasks in MujocoCar; (b) The RL with an Assistant Reward Agent (ReLara) algorithm Ma et al., 2024b , which learns an additional reward agent to generate informative rewards.
-ReLara can be considered as a special case of CenRA where each policy agent is paired with a separate reward agent, without cross-task information sharing; (c) The Policy Optimization and Policy Correction (PiCor) algorithm Bai et al., (2023), which uses policy correction to adapt and transfer information across policies among individual tasks; (d) The Multi-Critic Actor Learning (MCAL) algorithm Mysore et al., (2022), which maintains separate critics for each task for value estimation and knowledge representation, while training a unified multi-task actor.
-We use the CleanRL library Huang et al., (2022) or the author-provided codes to implement these baselines.
-(a) 2DMaze
-(b) 3DPickup
-(c) MujocoCar
-Figure 4: Comparison of the learning performance of CenRA with the baselines.
-Table 1: Comparison of episodic returns and standard errors.
-Algo. 2DMaze
-Maze #1 Maze #2 Maze #3 Maze #4
-CenRA 0.89 $\pm$ 0.03 0.91 $\pm$ 0.02 0.92 $\pm$ 0.02 0.93 $\pm$ 0.02
-DQN 0.77 $\pm$ 0.05 0.49 $\pm$ 0.09 0.54 $\pm$ 0.09 0.78 $\pm$ 0.05
-ReLara 0.83 $\pm$ 0.05 0.88 $\pm$ 0.03 0.76 $\pm$ 0.05 0.74 $\pm$ 0.05
-PiCor 0.80 $\pm$ 0.05 0.85 $\pm$ 0.05 0.78 $\pm$ 0.06 0.84 $\pm$ 0.05
-MCAL 0.76 $\pm$ 0.07 0.72 $\pm$ 0.09 0.57 $\pm$ 0.08 0.69 $\pm$ 0.08
-Algo. 3DPickup
-Ball Cube Key Health kit
-CenRA 0.95 $\pm$ 0.02 0.68 $\pm$ 0.09 0.80 $\pm$ 0.06 0.69 $\pm$ 0.07
-DQN 0.34 $\pm$ 0.06 0.07 $\pm$ 0.03 0.28 $\pm$ 0.03 0.28 $\pm$ 0.07
-ReLara 0.45 $\pm$ 0.08 0.39 $\pm$ 0.10 0.33 $\pm$ 0.09 0.29 $\pm$ 0.08
-PiCor 0.28 $\pm$ 0.07 0.44 $\pm$ 0.08 0.35 $\pm$ 0.09 0.28 $\pm$ 0.10
-MCAL 0.24 $\pm$ 0.07 0.34 $\pm$ 0.05 0.36 $\pm$ 0.08 0.45 $\pm$ 0.07
-Algo. MujocoCar
-Target #1 Target #2 Target #3 Target #4
-CenRA 588.2 $\pm$ 0.73 549.3 $\pm$ 0.64 447.7 $\pm$ 0.67 474.3 $\pm$ 0.66
-SAC 247.3 $\pm$ 0.53 210.6 $\pm$ 0.49 177.7 $\pm$ 0.41 156.4 $\pm$ 0.38
-ReLara 434.3 $\pm$ 0.68 436.0 $\pm$ 0.67 308.1 $\pm$ 0.58 420.8 $\pm$ 0.69
-PiCor 411.0 $\pm$ 0.66 426.0 $\pm$ 0.63 470.7 $\pm$ 0.66 442.5 $\pm$ 0.70
-MCAL 341.0 $\pm$ 0.62 315.6 $\pm$ 0.58 430.4 $\pm$ 0.59 389.8 $\pm$ 0.59
+We benchmark CenRA against several state-of-the-art baselines: (a) the backbone RL algorithms of the policy agents: DQN (Mnih et al.,, 2015) for discrete control tasks and SAC (Haarnoja et al., 2018b, ) for continuous control tasks; (b) the ReLara algorithm (Ma et al., 2024a, ), which can be regarded as a decentralized variant of CenRA, where each policy agent is paired with a separate reward agent, without cross-task information sharing; (c) the Policy Optimization and Policy Correction (PiCor) algorithm (Bai et al.,, 2023); (d) the Multi-Critic Actor Learning (MCAL) algorithm (Mysore et al.,, 2022); (e) the Parameter-compositional MTRL (PaCo) algorithm (Sun et al.,, 2022); and (f) the MTRL with Soft Modularization (SoftModule) (Yang et al.,, 2020).
+They are implemented by either the CleanRL library (Huang et al.,, 2022) or official codebases.
+Each task is trained with 10 different random seeds, and the average results are reported.
+In the Meta-World domain, ML10-sparse provides 10 training tasks and 5 held-out test tasks, while ML50-sparse includes 45 training tasks and 5 test tasks.
+For the remaining domains, each consists of 4 training tasks and 1 test task.
+In this section, we evaluate the final returns achieved by the trained agents, averaged over all training tasks in each domain, as shown in Table 1.
+We additionally report the episodic returns and their standard errors throughout training in the 2DMaze, 3DPickup, and MujocoCar domains in Figure 4.
+To ensure a fair comparison, we adopt consistent hyperparameters (where applicable) and identical network architectures across all experiments; detailed configurations are provided in Appendix B.
+Figure 4: Comparison of CenRA with baselines in 2DMaze, 3DPickup, and MujocoCar domains.
+Table 1: Episodic returns (mean $\pm$ standard error) of all trained agents tested over 100 episodes and averaged across all training tasks in each domain ($\uparrow$ higher is better).
+Algorithm ML10-sparse ML50-sparse 2DMaze 3DPickup MujocoCar
+CenRA (ours) 0.875 $\pm$ 0.121 0.755 $\pm$ 0.034 0.913 $\pm$ 0.023 0.880 $\pm$ 0.060 514.875 $\pm$ 0.675
+DQN/SAC 0.256 $\pm$ 0.056 0.189 $\pm$ 0.012 0.645 $\pm$ 0.070 0.243 $\pm$ 0.048 198.000 $\pm$ 0.453
+ReLara 0.674 $\pm$ 0.105 0.541 $\pm$ 0.057 0.803 $\pm$ 0.065 0.565 $\pm$ 0.088 429.800 $\pm$ 0.655
+PiCor 0.865 $\pm$ 0.230 0.672 $\pm$ 0.123 0.818 $\pm$ 0.053 0.438 $\pm$ 0.085 437.550 $\pm$ 0.663
+MCAL 0.842 $\pm$ 0.067 0.605 $\pm$ 0.055 0.885 $\pm$ 0.080 0.548 $\pm$ 0.068 369.200 $\pm$ 0.595
+PaCo 0.854 $\pm$ 0.045 0.582 $\pm$ 0.022 0.834 $\pm$ 0.057 0.557 $\pm$ 0.072 421.210 $\pm$ 0.635
+SoftModule 0.630 $\pm$ 0.042 0.423 $\pm$ 0.057 0.822 $\pm$ 0.076 0.486 $\pm$ 0.055 355.125 $\pm$ 0.594
+We observe that CenRA consistently outperforms all baselines in three main aspects.
+First, it achieves the highest episodic returns in all tasks, demonstrating superior learning efficiency and faster convergence.
+Moreover, it demonstrates good stability and robustness, exhibiting fewer fluctuations and oscillations, especially after convergence, compared to other models.
+Notably, all tasks provide only sparse rewards, CenRA addresses this challenge through the auxiliary dense rewards with meaningful information, effectively guiding learning.
+This mechanism not only distinguishes CenRA from other structurally shared methods, but also provides a targeted solution to the sparse-reward problem.
+Second, while baselines like PiCor and MCAL often show uneven progress across different tasks within the same domain, CenRA maintains well-balanced performance by showing relatively consistent learning progress and minimal variability across each four-task groups.
+This ensures that no single task dominates or lags behind, which is crucial in multi-task learning.
+Third, the CRA effectively enhances knowledge sharing among tasks.
+This is evident from the comparison with ReLara, which uses independent reward agents and lacks the mechanism for knowledge exchange.
+By extracting and distributing insights from one task to another, the CRA improves the learning efficiency of individual tasks, highlighting the advantages of integrated knowledge management.
+5.2 What Has the Centralized Reward Agent Learned?
+In this section, we visualize the learned knowledge rewards by the centralized reward agent $\mathcal{A}^{rwd}$ in the 2DMaze environment.
+After training on the four tasks in Section 5.1 of the paper, we let the CRA generate the knowledge rewards for each action in every state and visualize the action direction that yields the maximum rewards, $a^{*}=\operatorname*{arg\,max}_{a}{{\pi^{rwd}}^{*}(s_{i},a)},s_{i}\sim S$, in Figure 5.
 (a) Maze #1
@@
 Figure 5: Visualization of the action yielding the maximum knowledge rewards in the four 2DMaze tasks.
-We train multiple instances with different random seeds for each task to report the average data.
-The results of the episodic returns and their standard errors throughout training are shown in Figure 4 and Table 1.
-To ensure a fair comparison, we use the same hyperparameters (if applicable) and network structures across all experiments, the details are provided in Appendix B.
-We observe that CenRA consistently outperforms the baselines in three main aspects.
-First, CenRA achieves the highest episodic returns in nearly all tasks, demonstrating superior learning efficiency and faster convergence.
-Importantly, it also maintains good stability and robustness, which can be seen in the significantly fewer fluctuations and oscillations especially after convergence, compared to other models.
-Second, unlike PiCor and MCAL, which have shown uneven progress across different tasks in the same environment, CenRA maintains a well-balanced performance by showing relatively consistent learning progress and minimal variability in learning outcomes across each four-task group.
-This ensures that no single task’s performance is disproportionately high at the expense of others, which is crucial in multi-task learning.
-Third, the centralized reward agent in CenRA effectively enhances knowledge sharing among tasks.
-This is evident from the comparison with ReLara, which uses independent reward agents and lacks the mechanism for knowledge exchange.
-By extracting and distributing insights from one task to another, the CRA improves the learning efficiency of individual tasks, highlighting the advantages of integrated knowledge management.
-5.2 What Has the Centralized Reward Agent Learned?
-In this section, we visualize the learned knowledge rewards by the centralized reward agent $\mathcal{A}^{rwd}$ in the 2DMaze environment.
-After training on the four tasks in Section 5.1 of the paper, we let the CRA generate the knowledge rewards for each action in every state and visualize the action direction that yields the maximum rewards, $a^{*}=\operatorname*{arg\,max}_{a}{{\pi^{rwd}}^{*}(s_{i},a)},s_{i}\sim S$, in Figure 5.
 The shaded areas in the figures represent regions within the real task that the agent cannot reach, as it cannot access the space behind the door without picking up the key.
@@
 5.3 Knowledge Transfer to New Tasks
-In this section, we assess the centralized reward agent’s ability to transfer learned knowledge to new tasks.
-We continue using the CRA model trained on the four tasks in Section 5.1 but initiate a new policy agent to tackle a newly designed task that the models have never encountered before.
-For the CenRA framework, we explore two scenarios: one where the CRA continues to optimize while cooperating with the new policy agent (CenRA w/ learning), and another where only the policy agent is trained without further updates to the CRA (CenRA w/o learning).
-We compare the two approaches with the backbone algorithms (DQN and SAC) and ReLara.
-In the ReLara model, the reward agent starts training anew (without pre-learned knowledge).
-The results are shown in Figure 6 and Table 2.
-We observe that CenRA with further learning achieved rapid convergence, mainly due to the CRA’s retention of previously acquired knowledge and its ability to quickly adapt to new tasks through continuous optimization.
-Notably, even without any additional training, CenRA outperforms both the ReLara, which requires training a new reward agent, and the backbone algorithms, which lack additional information.
-In the CenRA framework, the CRA stores environment-relevant knowledge and transfers it to other policy agents, thus improving their learning efficiency.
-Additionally, sharing pre-learned information with policy agents is particularly crucial in our experiments with challenging sparse rewards.
+In this section, we assess the CRA’s ability to transfer previously learned knowledge to unseen tasks.
+Specifically, we continue using the trained CRA model in Section 5.1, while initializing new policy agents to tackle new tasks from the same domain.
+These include 5 test tasks for ML10-sparse and ML50-sparse, and 1 test task for each of the remaining domains, none of which were encountered during the initial training.
+For the CenRA, we explore two scenarios: (1) the CRA continues to be optimized in collaboration with the new policy agent (CenRA w/ learning); and (2) only the policy agent is updated while the CRA remains fixed, relying only on its previously acquired knowledge (CenRA w/o learning).
+We compare the two settings against the backbone algorithms and ReLara.
+In ReLara, the reward agent is trained anew without pre-learned knowledge.
+The results are presented in Figure 6 and Table 2.
+We observe that CenRA with further learning achieved rapid convergence, mainly due to the CRA’s ability to retain previously acquired knowledge while continuing to adapt to new tasks through ongoing optimization.
+Remarkably, even without any additional training, CenRA still outperforms both ReLara, which requires training a new reward agent, and the backbone algorithms, which lack additional information.
+This advantage stems from the CRA’s ability to encode and transfer environment-relevant knowledge, which can then be directly reused by new policy agents to guide their learning.
+Such knowledge transfer is particularly critical in our experiments involving challenging sparse-reward tasks.
 Without any external knowledge, learning would require extensive exploration.
@@
 Figure 6: Comparison of the learning performance of CenRA with the baselines in new tasks.
-Table 2: Comparison of the episodic returns and standard errors in new tasks.
-Algo. New 2DMaze New 3DPickup New MujocoCar
-CenRA w/ learning 0.95 $\pm$ 0.01 0.96 $\pm$ 0.00 532.00 $\pm$ 0.61
-CenRA w/o learning 0.89 $\pm$ 0.03 0.68 $\pm$ 0.00 524.73 $\pm$ 0.60
-ReLara 0.76 $\pm$ 0.06 0.26 $\pm$ 0.00 224.65 $\pm$ 0.49
-DQN/SAC 0.26 $\pm$ 0.08 0.16 $\pm$ 0.00 129.06 $\pm$ 0.30
+Table 2: Episodic returns (mean $\pm$ standard error) of all trained agents in the new tasks, tested over 100 episodes in each domain ($\uparrow$ higher is better).
+Algorithm ML10-sparse ML50-sparse 2DMaze 3DPickup MujocoCar
+CenRA w/ learning 0.902 $\pm$ 0.021 0.824 $\pm$ 0.012 0.952 $\pm$ 0.010 0.963 $\pm$ 0.002 532.080 $\pm$ 1.610
+CenRA w/o learning 0.887 $\pm$ 0.011 0.809 $\pm$ 0.009 0.894 $\pm$ 0.032 0.678 $\pm$ 0.003 524.727 $\pm$ 0.588
+ReLara 0.702 $\pm$ 0.086 0.612 $\pm$ 0.012 0.759 $\pm$ 0.056 0.263 $\pm$ 0.002 224.648 $\pm$ 0.492
+DQN/SAC 0.228 $\pm$ 0.105 0.210 $\pm$ 0.034 0.263 $\pm$ 0.084 0.158 $\pm$ 0.003 129.055 $\pm$ 0.296
 (a) Before getting the key.
@@
 Figure 7: Actions yielding the highest knowledge rewards at each state in a new 2DMaze task.
-To further demonstrate CenRA’s transferability, we select the 2DMaze environment to visualize the learned knowledge of CRA in an unseen task.
-As shown in Figure 7, we plot the directions of actions that yield the maximum knowledge reward at each position, categorized into two scenarios before and after obtaining the key.
-While some guidance in peripheral areas may be slightly off-target, most states receive reasonable rewards that align with human understanding.
-This demonstrates the effective transfer of knowledge and with such dense rewards, the agent’s adaptation to new tasks is able to be well supported.
+To further demonstrate CenRA’s transferability, we select the 2DMaze environment to visualize the knowledge provided by CRA when facing an unseen task.
+As shown in Figure 7, we plot the directions of actions that yield the maximum knowledge reward at each position, categorized into two scenarios: before and after obtaining the key.
+While some guidance in peripheral regions may appear slightly misaligned, most states receive reasonable rewards that align with human understanding.
+This demonstrates the effectiveness of knowledge transfer, and with such dense rewards, the agent’s adaptation to new tasks is able to be well supported.
 5.4 Effect of Sampling Weight
-We conducted experiments to understand the effects of the information synchronization mechanism in the CenRA framework.
+We conducted experiments to understand the effects of the information synchronization mechanism in the CenRA.
 Specifically, we compare the full CenRA model against three variants: (a) CenRA without the similarity weight $\bm{w}^{sim}$, (b) CenRA without the performance weight $\bm{w}^{per}$, and (c) CenRA without the entire sampling weight.
-The comparison results are shown in Figure 8 and Table 3.
-(a) 2DMaze
-(b) 3DPickup
-(c) MujocoCar
-Figure 8: Comparison of CenRA with ablation of different batch sampling control weights.
+To better illustrate the differences among tasks and highlight the role of sampling weights in task coordination and synchronization, we select the four-task domains, i.e., 2DMaze, 3DPickup, and MujocoCar.
+The comparison results are shown in Table 3.
 Table 3: Comparison of CenRA with ablation of different batch sampling control weights.
@@
 While the full CenRA model does not always achieve the lowest variance, it consistently outperforms the other three ablation models regarding overall system performance.
-Both weights are essential for the information synchronization mechanism, with the performance weight $\bm{w}^{per}$ having a more important effect.
-This weight enables CRA to pay more attention to the policy agents that are underperforming or progressing slowly.
-This suggests that considering the overall learning performance of the multi-task system is also an important objective CenRA seeks to achieve.
+Both weights play essential roles in the information synchronization, with the performance weight $\bm{w}^{per}$ having a more important impact.
+It allows CRA to pay more attention to the policy agents that are underperforming or progressing slowly.
+This suggests that considering the overall learning performance of the multi-task system is also an important objective that CenRA seeks to achieve.
 6 Discussion and Conclusion
 We propose a novel framework named CenRA that innovatively integrates reward shaping into MTRL.
-The framework not only can share environment-relevant knowledge across tasks to improve learning efficiency, but also effectively addresses the sparse-reward challenge.
-Specifically, a CRA functions as a knowledge pool, responsible for knowledge distillation and distribution across multiple tasks.
+The framework not only shares domain knowledge across tasks to improve learning efficiency but also effectively addresses the sparse-reward challenge.
+Specifically, the centralized reward agent functions as a knowledge pool, responsible for knowledge distillation and distribution across multiple tasks.
 Furthermore, our information synchronization mechanism mitigates knowledge distribution disparities across tasks, ensuring optimal system-wide performance.
 Experiments demonstrate that the dense knowledge rewards generated by the CRA efficiently guide the policy agents’ learning, leading to faster convergence compared to baseline methods.
-The CenRA framework proves to be robust and performs superiorly in both multi-task learning and new tasks adaptation.
-CenRA’s current limitation is its requirement for uniform state and action spaces across tasks.
-Future improvements could explore preprocessing techniques to adapt the framework for varying state and action spaces to broaden CenRA’s applicability.
+The CenRA also performs superiorly and robustly in transferring acquired knowledge to new tasks.
+CenRA’s current limitation is its requirement for uniform state and action shapes across tasks.
+Future improvements could explore preprocessing techniques to adapt the framework for varying state and action shapes to broaden CenRA’s applicability.
 Additionally, the fixed trade-off between similarity weight and performance weight may not be ideal.
-A more flexible approach, potentially through adaptive regulation of these weights, could enhance the framework further.
-Besides, the performance weight might favor underperforming tasks, achieving overall balance but potentially limiting the upper bound of top-performing tasks.
-This issue could necessitate a better mechanism for trade-off.
-7 Acknowledgement
+A more flexible approach, potentially through adaptive regulation of these weights, could further enhance the framework.
+Besides, the performance weight might favor underperforming tasks, achieving overall balance, but potentially limiting the upper bound of top-performing tasks.
+This issue could necessitate a better mechanism for trade-offs.
 Appendix A Mutli-Task Experimental Configurations
-In Section 5 Experiments, we conduct experiments in three environments with multiple tasks: 2DMaze, 3DPickup, and MujocoCar. The detailed configurations of each task are illustrated in Figure 9.
-Figure 9: The configurations of multiple tasks used in your experiments.
-Appendix B Network Structures and Hyperparameters
+We conduct experiments in four domains with multiple tasks: Meta-World, 2DMaze, 3DPickup, and MujocoCar.
+The detailed configurations of each task are illustrated in Figure 8.
+The Meta-World tasks illustration is adapted from (Yu et al., 2020b, ).
+Figure 8: Illustration of multiple tasks in different domains in our experiments.
+Appendix B Experimental Configurations
 B.1 Network Structures
@@
 Figure 9: The structures of neural networks in our implementation.
-B.2 Hyperparameters
+B.2 Computing Resources
+The experiments in this paper were conducted on a computing cluster, with the detailed hardware configurations listed in Table 4.
+Table 4: The computing resources used in the experiments.
+Component Specification
+Operating System (OS) Ubuntu 20.04
+Central Processing Unit (CPU) 2x Intel Xeon Gold 6326
+Random Access Memory (RAM) 256GB
+Graphics Processing Unit (GPU) 1x NVIDIA A100 20GB
+Brand Supermicro 2022
+B.3 Hyperparameters
 We have observed that CenRA demonstrated high robustness and was not sensitive to hyperparameter choices. Table 5 shows the hyperparameters we used in all the experiments.

### v3

@@
 Abstract
-Reward shaping is effective in addressing the sparse-reward challenge in reinforcement learning by providing immediate feedback through auxiliary informative rewards.
+Reward shaping is effective in addressing the sparse-reward challenge in reinforcement learning (RL) by providing immediate feedback through auxiliary, informative rewards.
 Based on the reward shaping strategy, we propose a novel multi-task reinforcement learning framework that integrates a centralized reward agent (CRA) and multiple distributed policy agents.
-The CRA functions as a knowledge pool, which aims to distill knowledge from various tasks and distribute it to individual policy agents to improve learning efficiency.
-Specifically, the shaped rewards serve as a straightforward metric to encode knowledge.
+The CRA functions as a knowledge pool, aimed at distilling knowledge from various tasks and distributing it to individual policy agents to improve learning efficiency.
+Specifically, the shaped rewards serve as a straightforward metric for encoding knowledge.
 This framework not only enhances knowledge sharing across established tasks but also adapts to new tasks by transferring meaningful reward signals.
-We validate the proposed method on both discrete and continuous domains, including the representative meta world benchmark, demonstrating its robustness in multi-task sparse-reward settings and its effective transferability to unseen tasks.
+We validate the proposed method on both discrete and continuous domains, including the representative Meta-World benchmark, demonstrating its robustness in multi-task sparse-reward settings and its effective transferability to unseen tasks.
 1 Introduction
-Reinforcement learning (RL) has made significant progress across various domains, like robotics (Kober et al.,, 2013), gaming (Lample and Chaplot,, 2017), autonomous vehicles (Aradi,, 2020), signal processing (Luo et al.,, 2024), and large language models (Shinn et al.,, 2023; Ouyang et al.,, 2022).
-However, environments with sparse and delayed rewards remain a significant challenge, as the absence of immediate feedback hinders the agent from distinguishing the value of states and leads to aimless exploration (Ladosz et al.,, 2022).
-Reward Shaping (RS) has been proven to be an effective technique to address this challenge by providing additional dense and informative rewards (Sorg et al., 2010b, ; Sorg et al., 2010a, ).
-Concurrently, multi-task reinforcement learning (MTRL) is becoming increasingly important for its ability to share and transfer knowledge across tasks.
-In this context, the auxiliary rewards infused with task-specific information in RS offer a straightforward entry to distribute knowledge among different tasks.
-Therefore, we believe that integrating RS techniques into MTRL is a highly promising and intuitive direction to enhance the efficacy of multi-task system learning.
-Presently, numerous MTRL algorithms involving knowledge transfer have been developed.
-Policy distillation methods identify and combine the commonalities across different policies (Rusu et al.,, 2016; Teh et al.,, 2017; Parisotto et al.,, 2016; Xu et al.,, 2024; Ma et al., 2024b, ; Ma et al.,, 2023); representation sharing methods extract and disseminate the common features or gradients among agents (Yang et al.,, 2020; D’Eramo et al.,, 2020; Sodhani et al.,, 2021); and parameter sharing methods design architectural modules to reuse parameters or layers across networks (Sun et al.,, 2022; Cheng et al.,, 2023).
-Despite their potential, these strategies often encounter challenges such as slow adaptation and response to the transferred knowledge, or suffer from delays in reutilizing and comprehending this information.
-Therefore, leveraging the RS mechanism, which directly adds a metric to the reward function, offers a compelling alternative to address these limitations.
-Regarding the RS techniques, not all shaped rewards serve effectively as the medium for knowledge transfer.
-Specifically, the intrinsic motivation based rewards are typically designed using heuristics to generate task-agnostic signals.
-Examples include incorporating exploration bonuses (Bellemare et al.,, 2016; Ostrovski et al.,, 2017; Devidze et al.,, 2022), rewarding novel states (Tang et al.,, 2017; Burda et al.,, 2018), and encouraging curiosity-driven behaviors (Pathak et al.,, 2017; Mavor-Parker et al.,, 2022).
-Although these approaches broaden extensive exploration, they are not directly related to specific tasks and thus lack transferability.
-Consequently, we focus on another branch of RS methods, the task-contextual rewards, which automatically learn and encode task-specific information, such as the hidden values, contributions of states, or future-oriented insights, that can be shared effectively across various tasks (Ma et al., 2024a, ; Ma et al., 2025b, ; Ma et al., 2025a, ; Mguni et al.,, 2023; Memarian et al.,, 2021).
-To share task-related knowledge in MTRL via RS techniques, and inspired by the ReLara framework (Ma et al., 2024a, ), which integrates an assistant reward agent to densify sparse environmental rewards, we propose the Centralized Reward Agent based MTRL fRAmework (CenRA).
+Reinforcement learning (RL) has made significant progress across various domains, such as robotics [Kober et al., 2013], gaming [Lample and Chaplot, 2017], autonomous vehicles [Aradi, 2020], signal processing [Luo et al., 2024], and large language models [Shinn et al., 2023, Ouyang et al., 2022].
+However, environments with sparse and delayed rewards remain a significant challenge, as the absence of immediate feedback hinders the agent from distinguishing the value of states and leads to aimless exploration [Ladosz et al., 2022].
+Reward Shaping (RS) has been proven to be an effective technique for addressing this challenge by providing additional dense and informative rewards [Sorg et al., 2010b, a].
+Concurrently, multi-task reinforcement learning (MTRL) is becoming increasingly important due to its ability to transfer knowledge across tasks.
+In this context, the auxiliary rewards infused with task-specific information in RS offer a straightforward means to distribute knowledge among different tasks.
+Integrating RS techniques into MTRL is a highly promising and intuitive direction to enhance the efficacy of multi-task learning systems.
+Numerous MTRL algorithms for knowledge transfer have been developed.
+Policy distillation methods identify and combine commonalities across different policies [Rusu et al., 2016, Teh et al., 2017, Parisotto et al., 2016, Xu et al., 2024]; representation sharing methods extract and share the common features or gradients among agents [Yang et al., 2020, D’Eramo et al., 2020, Sodhani et al., 2021]; and parameter sharing methods design architectural modules to reuse parameters or layers across networks [Sun et al., 2022, Cheng et al., 2023].
+Despite their potential, these strategies often face slow adaptation to and limited utilization of transferred knowledge.
+Therefore, leveraging reward shaping, which directly adds a metric to the reward function, offers a compelling alternative to address these limitations.
+Regarding reward shaping, not all shaped rewards effectively serve as a medium for knowledge transfer.
+Specifically, the intrinsic-motivation-based rewards are typically designed using heuristics to generate task-agnostic signals.
+Examples include incorporating exploration bonuses [Bellemare et al., 2016, Ostrovski et al., 2017, Devidze et al., 2022], rewarding novel states [Tang et al., 2017, Burda et al., 2018], and encouraging curiosity-driven behaviors [Pathak et al., 2017, Mavor-Parker et al., 2022].
+Although these approaches encourage broader exploration, they are not directly related to specific tasks and thus lack transferability.
+Consequently, we focus on another branch of RS methods, task-contextual rewards, which automatically learn and encode task-specific information, such as hidden values, states contributions, or future-oriented insights, that can be effectively shared across various tasks [Ma et al., 2024a, 2025b, 2025a, Mguni et al., 2023, Memarian et al., 2021].
+To share task-related knowledge in MTRL via RS techniques, and inspired by the ReLara framework [Ma et al., 2024a], which integrates an assistant reward agent to densify sparse environmental rewards, we propose the Centralized Reward Agent based MTRL fRAmework (CenRA) [^1] The source code is accessible at: https://github.com/mahaozhe/CenRA.
 The framework consists of two main components: a centralized reward agent (CRA) and multiple distributed policy agents.
-Each policy agent individually learns control behaviors within their respective tasks and shares their experiences with the CRA.
+Each policy agent individually learns control behaviors within its respective tasks and shares its experiences with the CRA.
 The CRA extracts common knowledge from these experiences and learns to generate dense rewards that are encoded with task-specific information.
 These rewards are then distributed back to the policy agents to augment their original environmental rewards.
-Additionally, given that different tasks may contribute variably to the MTRL system, we introduce an information synchronization mechanism to further balance the knowledge distribution, considering task similarity and agent learning progress, to ensure a system-wide optimal performance.
+Additionally, given that different tasks may contribute variably to the MTRL system, we introduce an information synchronization mechanism to further balance knowledge distribution by considering task similarity and agent learning progress, thereby ensuring system-wide optimal performance.
 The main contributions of this paper are summarized as follows:
   1. (i)
-    We propose the CenRA framework targeted to solve MTRL problems.
-It incorporates a CRA that functions as a centralized knowledge pool, efficiently distilling and distributing valuable information from various tasks to policy agents, and further adapting to new tasks.
+    We propose the CenRA framework to address MTRL problems.
+It incorporates a CRA that functions as a knowledge pool, efficiently distilling and distributing valuable information from various tasks to policy agents while adapting to new tasks.
   2. (ii)
-    CenRA leverages the advantage of reward shaping techniques to infuse insights via dense rewards.
-This approach not only provides a direct metric for policy agents to absorb knowledge but also effectively addresses the sparse-reward challenge.
+    CenRA leverages reward shaping techniques to infuse insights via dense rewards.
+This approach not only provides a direct signal for policy agents to absorb knowledge but also effectively addresses the sparse-reward challenge.
   3. (iii)
-    We introduce an information synchronization mechanism that considers both task similarity and agent convergence progress to balance multi-task learning.
+    We introduce an information synchronization mechanism that considers both task similarity and agent learning progress to balance multi-task learning.
 This mechanism provides a novel direction for maintaining system equilibrium in MTRL.
   4. (iv)
-    CenRA is validated in both discrete and continuous control MTRL environments, where the extrinsic rewards are sparse.
-CenRA outperforms several baseline models in terms of learning efficiency, knowledge transferability, and system-wide performance.
+    CenRA is validated in both discrete and continuous control MTRL environments with sparse extrinsic rewards.
+CenRA outperforms baseline models in learning efficiency, knowledge transferability, and system-wide performance.
 2 Related Work
-Multi-task reinforcement learning (MTRL) has attracted significant attention recently due to its potential to share knowledge across multiple tasks, thereby improving learning performance (Caruana,, 1993).
+Multi-task reinforcement learning (MTRL) has attracted significant attention recently due to its potential to share knowledge across multiple tasks, thereby improving learning performance [Caruana, 1993].
 We discuss existing MTRL literature from three main directions:
-Knowledge Transfer methods focus on identifying and transferring task-relevant features across diverse tasks (Zeng et al.,, 2021).
-Policy distillation (Rusu et al.,, 2016) is a well-studied approach to extract and share task-specific behaviors or representations that many works are built on: Teh et al., (2017) introduced Distral, which distills a centroid policy from multiple task-policies; Parisotto et al., (2016) developed Actor-Mimic, where a single policy is trained to mimic several expert policies from different tasks; while Yin and Pan, (2017) incorporated hierarchical prioritized experience replay buffer to select and learn multi-task experiences; Hessel et al., (2019) further proposed an adaptation mechanism to equalize the impact of each task in policy distillation.
-Additionally, Xu et al., (2020) explored the transfer of offline knowledge to train policies, and further leveraged online learning for fine-tuning.
-Bai et al., (2023) introduced a dual-phase learning approach, optimizing individual policies while correcting them across multiple tasks.
-Mysore et al., (2022) used separate critics for each task to accompany a single actor to integrate their feedback.
+Knowledge Transfer methods focus on identifying and transferring task-relevant features across diverse tasks [Zeng et al., 2021].
+Policy distillation [Rusu et al., 2016] is a well-studied approach to extract and share task-specific behaviors or representations that many works are built on: Teh et al.
+[2017] introduced Distral, which distills a centroid policy from multiple task-policies; Parisotto et al.
+[2016] developed Actor-Mimic, where a single policy is trained to mimic several expert policies from different tasks; while Yin and Pan [2017] incorporated hierarchical prioritized experience replay buffer to select and learn multi-task experiences; Hessel et al.
+[2019] further proposed an adaptation mechanism to equalize the impact of each task in policy distillation.
+Additionally, Xu et al.
+[2020] explored the transfer of offline knowledge to train policies, and further leveraged online learning for fine-tuning.
+Bai et al.
+[2023] introduced a dual-phase learning approach, optimizing individual policies while correcting them across multiple tasks.
+Mysore et al.
+[2022] used separate critics for each task to accompany a single actor to integrate their feedback.
 These methods mitigate gradient interference to an extent, however, balancing the distribution of knowledge across tasks is crucial.
 Without a careful trade-off, the performance of the entire system could be compromised.
-Representation Sharing methods explore architectural solutions of reusing network modules or representing commonalities to the MTRL problem (D’Eramo et al.,, 2020; Devin et al.,, 2017; Hong et al.,, 2021).
-Sun et al., (2022) used a parameter compositional approach to learn and share a subspace of parameters, allowing policies for various tasks to be interpolated within it.
-Yang et al., (2020) employed soft modularization to learn foundational policies and utilized a routing network to generate probabilities to combine them.
-He et al., (2024) introduced the Dynamic Depth Routing framework, which dynamically adjusts the use of network modules in response to task difficulty.
-Sodhani et al., (2021) leveraged task-related metadata to create composable representations.
-Cheng et al., (2023) and Lan et al., (2023) both incorporated attention mechanisms: the former employed attention-based mixture of experts to capture task relationships, while the latter used Temporal Attention for contrastive learning purposes.
+Representation Sharing methods explore architectural solutions of reusing network modules or representing commonalities to the MTRL problem [D’Eramo et al., 2020, Devin et al., 2017, Hong et al., 2021, Ma et al., 2024b, 2023].
+Sun et al.
+[2022] used a parameter compositional approach to learn and share a subspace of parameters, allowing policies for various tasks to be interpolated within it.
+Yang et al.
+[2020] employed soft modularization to learn foundational policies and utilized a routing network to generate probabilities to combine them.
+He et al.
+[2024] introduced the Dynamic Depth Routing framework, which dynamically adjusts the use of network modules in response to task difficulty.
+Sodhani et al.
+[2021] leveraged task-related metadata to create composable representations.
+Cheng et al.
+[2023] and Lan et al.
+[2023] both incorporated attention mechanisms: the former employed attention-based mixture of experts to capture task relationships, while the latter used Temporal Attention for contrastive learning purposes.
 Although these methods demonstrate efficacy in learning shared representations, they may struggle to fully capture the complexity of highly diverse tasks.
@@
 Single-Policy Generalization methods learn a single policy to solve multiple tasks simultaneously or continuously, in the absence of information from prior policies or task-specific details, in which case, the primary goal is to enhance the policy’s generalization capabilities.
-Model-free meta-learning techniques have been proposed to enhance the multi-task generalization (Finn et al.,, 2017).
-Yang et al., (2017) designed a sharing network structure that allows an agent to learn multiple tasks concurrently.
-Vuong et al., (2019) introduced a confidence-sharing agent to detect and define shared regions between tasks to support single policy learning.
-Wan et al., (2020) proposed a transfer learning framework to handle mismatches in state and action spaces.
-Additionally, several methods focus on overcoming gradient interference to enhance the generalization in various tasks (Chen et al.,, 2018; Yu et al., 2020a, ), while Ammar et al., (2014) developed a consecutive learning policy gradient approach.
+Model-free meta-learning techniques have been proposed to enhance the multi-task generalization [Finn et al., 2017].
+Yang et al.
+[2017] designed a sharing network structure that allows an agent to learn multiple tasks concurrently.
+Vuong et al.
+[2019] introduced a confidence-sharing agent to detect and define shared regions between tasks to support single policy learning.
+Wan et al.
+[2020] proposed a transfer learning framework to handle mismatches in state and action spaces.
+Additionally, several methods focus on overcoming gradient interference to enhance the generalization in various tasks [Chen et al., 2018, Yu et al., 2020a], while Ammar et al.
+[2014] developed a consecutive learning policy gradient approach.
 These methods are efficient in saving computational resources, but the generalization ability of the policy may be constrained when faced with out-of-distribution or previously unseen tasks.
@@
 An MTRL agent aims to learn optimal policies $\pi_{i}:S\rightarrow A$ for each task $T_{i}\sim\mathcal{T}$, to maximize their corresponding expected cumulative rewards, or returns, denoted by $G_{i}=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}R_{i}(s_{t},a_{t})]$.
-RL with an Assistant Reward Agent (ReLara) (Ma et al., 2024a, ) introduces a dual-agent framework designed to tackle the challenge of sparse rewards in RL.
-The operational architecture of ReLara is shown in Figure 1.
+RL with an Assistant Reward Agent (ReLara) [Ma et al., 2024a] introduces a dual-agent framework designed to tackle the challenge of sparse rewards in RL.
 Within this framework, the original agent is termed as policy agent, while an assistant reward agent is integrated to enrich the feedback mechanism by generating dense, informative rewards.
-The reward agent, trained as a self-contained RL agent, autonomously extracts hidden value information from the environmental states (and potentially the actions of the policy agent) to craft meaningful reward signals.
+The reward agent, trained as a self-contained RL agent, autonomously extracts hidden value information from the environmental states and the actions of the policy agent to craft meaningful reward signals.
 These signals significantly improve learning efficiency by providing immediate and pertinent bonuses.
-Figure 1: The ReLara framework Ma et al., 2024a .
 4 Methodology
-We propose Centralized Reward Agent fRAmework (CenRA) for MTRL, which incorporates a centralized reward agent (CRA) to support simultaneous reinforcement learning agents across multiple tasks.
+We propose the Centralized Reward Agent fRAmework (CenRA) for MTRL, which incorporates a centralized reward agent (CRA) to support multiple reinforcement learning agents across multiple tasks.
 A high-level illustration of the CenRA framework is shown in Figure 1.
 The CRA is responsible for extracting general task-specific knowledge from various tasks and distributing valuable information to the policy agents by reconstructing their reward models.
-The detailed methodology for knowledge extraction and sharing is introduced in Section 4.1.
-Furthermore, to mitigate the potential disparities in the information that each task could contribute, which might lead to an imbalance in knowledge distribution, we introduce an information synchronization mechanism by considering two main factors: the similarity of the tasks and the online learning performance of the policy agents, details given in Section 4.2.
+The detailed methodology for knowledge extraction and sharing is presented in Section 4.1.
+Furthermore, to mitigate the potential disparities in the information that each task contributes, which might lead to an imbalance in knowledge distribution, we introduce an information synchronization mechanism by considering two main factors: the similarity of the tasks and the online learning performance of the policy agents, details given in Section 4.2.
 Finally, the overall framework of CenRA is presented in Section 4.3.
 Figure 1: A high-level illustration of the CenRA framework.
-4.1 Knowledge Extraction and Share
+The centralized reward agent functions as a knowledge repository, distilling information from various tasks and distributing it to individual policy agents to enhance learning efficiency.
+4.1 Knowledge Distillation and Distribution
 4.1.1 Problem Formulation
@@
 Each policy agent $\mathcal{A}^{pol}_{i}$ operates independently to complete its corresponding task $T_{i}$, utilizing appropriate RL algorithms as backbones.
-For example, implementing DQN (Mnih et al.,, 2015) for discrete control tasks, while TD3 (Fujimoto et al.,, 2018) or SAC (Haarnoja et al., 2018a, ) for continuous control tasks.
-Moreover, the policy of CRA $\mathcal{A}^{rwd}$ is $\pi^{rwd}$, and the internal policy of policy agent $\mathcal{A}^{pol}_{i}$ is $\pi^{pol}_{i}$, respectively.
+For example, implementing DQN [Mnih et al., 2015] for discrete control tasks, while TD3 [Fujimoto et al., 2018] or SAC [Haarnoja et al., 2018a] for continuous control tasks.
+Moreover, the policy of CRA $\mathcal{A}^{rwd}$ is $\pi^{rwd}$, and the internal policy of policy agent $\mathcal{A}^{pol}_{i}$ is $\pi^{pol}_{i}$.
 4.1.2 Centralized Reward Agent
 The CRA $\mathcal{A}^{rwd}$ aims to extract environment-relevant knowledge and distribute it to policy agents by generating additional dense rewards to support their original reward functions.
-Similar to the ReLara framework Ma et al., 2024a , we model the CRA as a self-contained RL agent, yet, as an extension to ReLara, our CRA is designed to concurrently interact with multiple policy agents and their respective tasks.
+Similar to the ReLara framework [Ma et al., 2024a], we model the CRA as a self-contained RL agent, yet, as an extension to ReLara, our CRA is designed to concurrently interact with multiple policy agents and their respective tasks.
 The CRA’s policy $\pi^{rwd}$ generates continuous rewards given both an environmental state and a policy agent’s behavior.
 Specifically, $\pi^{rwd}$ maps the Cartesian product of the state space and action space, $S\times A$, to a defined reward space, which constrains the rewards to a range of real numbers, $\mathcal{R}=[R_{min},R_{max}]\subset\mathbb{R}$.
-For simplicity’s sake, we denote the observation of CRA as $s^{rwd}=(s_{i},a_{i})$, where $s_{i}\sim T_{i}$ and $a_{i}\sim\pi^{pol}_{i}(s_{i})$.
+For simplicity, we denote the observation of the CRA as $s^{rwd}=(s_{i},a_{i})$, where $s_{i}\sim T_{i}$ and $a_{i}\sim\pi^{pol}_{i}(s_{i})$.
 To distinguish from the environmental reward, the generated reward is termed as knowledge reward, denoted as $r^{knw}$.
-We adopt an off-policy actor-critic algorithm to optimize the CRA Konda and Tsitsiklis, (1999).
+We adopt an off-policy actor-critic algorithm to optimize the CRA [Konda and Tsitsiklis, 1999].
 To aggregate and reuse experiences from all policy agents, a concatenated replay buffer $\mathcal{D}=\bigcup_{i=1}^{N}{\mathcal{D}_{i}}$ is constructed, where $\mathcal{D}_{i}$ represents the replay buffer of each policy agent $\mathcal{A}^{pol}_{i}$.
@@
 4.1.3 Policy Agents with Knowledge Rewards
-Each policy agent $\mathcal{A}^{pol}_{i}$ interacts with its specific environment and stores the experiences in its corresponding replay buffer $\mathcal{D}_{i}$.
-Policy agents receive two types of rewards: the environmental reward $r^{env}_{i}$ from their respective task $T_{i}$ and the knowledge reward $r^{knw}$ from CRA.
+Each policy agent $\mathcal{A}^{pol}_{i}$ stores the experiences in its corresponding replay buffer $\mathcal{D}_{i}$.
+They receive two types of rewards: the environmental reward $r^{env}_{i}$ from their respective task $T_{i}$ and the knowledge reward $r^{knw}$ from CRA.
 The augmented reward is given by:
@@
 4.2 Information Synchronization of Policy Agents
-In the CenRA framework, the information provided by different tasks may exhibit significant disparities, potentially leading to an imbalance in knowledge extraction and distribution.
-In this section, we introduce an information synchronization mechanism for CenRA to maintain a balanced manner from the perspective of the entire system.
-Specifically, this mechanism is implemented by controlling the quantity of samples that CRA retrieves from each task’s replay buffer $\mathcal{D}_{i}$ by a sampling weight $\bm{w}$.
-We mainly consider two aspects: the similarity among tasks and the real-time learning performance of the policy agents.
-Similarity Weight is derived from the similarity among tasks, enabling the CRA to focus on relatively outlier tasks.
+In the CenRA, the information provided by different tasks may exhibit significant disparities, potentially leading to an imbalance in knowledge extraction and distribution.
+We introduce an information synchronization mechanism for CenRA to maintain a balanced manner from the perspective of the entire system.
+Specifically, we control the quantity of samples that CRA retrieves from each task’s replay buffer $\mathcal{D}_{i}$ by a sampling weight $\bm{w}$, by considering two aspects: the similarity among tasks and the real-time learning performance of the policy agents.
+Similarity Weight $\bm{w}^{sim}$ is derived from the similarity among tasks, enabling the CRA to focus on relatively outlier tasks.
 To simplify computation, we use the hidden layers extracted from each policy agent’s neural network encoders to represent the tasks’ features.
 To reduce randomness, we average the hidden features of the most recent $K$ steps.
-We adopt a cross-attention mechanism, which is widely used in neural networks, to calculate the similarity weight Vaswani et al., (2017).
+We adopt a cross-attention mechanism to calculate the similarity weight [Vaswani et al., 2017].
 Specifically, for task $T_{i}$, let $\bm{H}_{i}$ denote the averaged hidden feature vector, which serves as the key, and the centroid of all tasks $\bm{c}$ acts as the query.
@@
 A larger $s_{i}$ indicates a greater similarity between $T_{i}$ and the centroid.
+It is worth noting that, to avoid the centroid $\bm{c}$ approaching zero due to feature vectors $\bm{H}_{i}$ pointing in opposite directions, all latent representations $\bm{H}_{i}$ in our framework are extracted from ReLU activation layers.
+This ensures that every element of $\bm{H}_{i}$ is non-negative, effectively preventing feature cancellation and maintaining a well-defined and numerically stable centroid $\bm{c}$.
 Given our assumption is that the tasks farther from the centroid require more attention, the similarity weight is defined as $\bm{w}^{sim}=\text{Softmax}\big([1/s_{1},1/s_{2},\dots,1/s_{N}]\big)$.
@@
 The performance weight is then defined as $\bm{w}^{per}=\text{Softmax}\big([1/R^{tail}_{1},1/R^{tail}_{2},\dots,1/R^{tail}_{N}]\big)$.
-Lastly, the final sampling weight $\bm{w}$ combines the similarity weight and the performance weight as $\bm{w}=\alpha\bm{w}^{sim}+(1-\alpha)\bm{w}^{per}$, where $\alpha$ is a hyperparameter to balance the two aspects.
+The final sampling weight $\bm{w}$ is formulated as $\bm{w}=\alpha\bm{w}^{sim}+(1-\alpha)\bm{w}^{per}$, where $\alpha$ is a hyperparameter to balance the two aspects.
 The CRA samples from each replay buffer $\mathcal{D}_{i}$ according to $\bm{w}$, ensuring a balanced and effective knowledge extraction and learning.
@@
 Sampling weights are calculated in real-time, using the most recently optimized encoders and the current learning performance, ensuring CRA continuously adjusts its focus to optimally balance knowledge extraction across multiple tasks.
-The learned CRA acts as a robust knowledge pool, which is able to support new tasks as an additional module by transferring knowledge through auxiliary reward signals.
-This is particularly beneficial in sparse-reward environments, as the knowledge rewards can guide the policy agents toward the correct direction and reduce the exploration burden.
+The learned CRA acts as a robust knowledge pool, which is able to support new tasks by transferring knowledge through auxiliary reward signals.
+This is particularly beneficial in sparse-reward environments, as the knowledge rewards can guide the policy agents toward the correct direction and reduce exploration burden.
 Additionally, the CRA can be further optimized alongside new tasks in a continuous learning scheme that enhances adaptability and effectiveness in dynamic settings.
@@
 2:  for each task $T_{i}$ do
-3:        $(s_{t},a_{t},r^{env}_{t},s_{t+1},a_{t+1})\sim\text{Interact}(\mathcal{A}^{pol}_{i},T_{i})$triangleright Interact and collect one transition
-4:        $r^{knw}_{t}\sim\mathcal{A}^{rwd}(s_{t},a_{t})$triangleright Sample an off-policy knowledge reward
+3:   $(s_{t},a_{t},r^{env}_{t},s_{t+1},a_{t+1})\sim\text{Interact}(\mathcal{A}^{pol}_{i},T_{i})$ $\triangleright$ Interact and collect one transition
+4:   $r^{knw}_{t}\sim\mathcal{A}^{rwd}(s_{t},a_{t})$ $\triangleright$ Sample an off-policy knowledge reward
 5:   $s^{rwd}_{t}=(s_{t},a_{t})$, $s^{rwd}_{t+1}=(s_{t+1},a_{t+1})$
-6:        $\mathcal{D}_{i}\leftarrow\mathcal{D}_{i}\cup\{(s^{rwd}_{t},r^{knw}_{t},r^{env}_{t},s^{rwd}_{t+1}|T_{i})\}$triangleright Store the transition in corresponding $\mathcal{D}_{i}$
-7:        Update policy agent $\mathcal{A}^{pol}_{i}$triangleright Update $\mathcal{A}^{pol}_{i}$ using backbone RL algorithm
+6:   $\mathcal{D}_{i}\leftarrow\mathcal{D}_{i}\cup\{(s^{rwd}_{t},r^{knw}_{t},r^{env}_{t},s^{rwd}_{t+1}|T_{i})\}$ $\triangleright$ Store the transition in corresponding $\mathcal{D}_{i}$
+7:   Update policy agent $\mathcal{A}^{pol}_{i}$ $\triangleright$ Update $\mathcal{A}^{pol}_{i}$ using backbone RL algorithm
 8:  end for
-9:     $\bm{w}=\alpha\bm{w}^{sim}+(1-\alpha)\bm{w}^{per}$triangleright Calculate sampling weight
-10:     $\{s^{rwd}_{t},r^{knw}_{t},r^{env}_{t},s^{rwd}_{t+1}|T_{i}\}_{\mathcal{B}}\sim\mathcal{D}|\bm{w}$triangleright Draw samples based on the sampling weight
+9:  $\bm{w}=\alpha\bm{w}^{sim}+(1-\alpha)\bm{w}^{per}$ $\triangleright$ Calculate sampling weight
+10:  $\{s^{rwd}_{t},r^{knw}_{t},r^{env}_{t},s^{rwd}_{t+1}|T_{i}\}_{\mathcal{B}}\sim\mathcal{D}|\bm{w}$ $\triangleright$ Draw samples based on the sampling weight
 11:  Update centralized reward agent $\mathcal{A}^{rwd}$
@@
 5 Experiments
-We conduct experiments in four MTRL domains as shown in Figure 3: the widely used Meta-World benchmark (including ML10 with 10 tasks and ML50 with 50 tasks) (Yu et al., 2020b, ), 2DMaze, 3DPickup (Chevalier-Boisvert et al.,, 2024), and MujocoCar (Ji et al.,, 2023).
-All tasks, including those in Meta-World, are crafted to provide sparse environmental rewards, where the agent receives a reward of $1$ only upon successful completion of the final objective, and 00 otherwise.
+We conduct experiments in four MTRL domains as shown in Figure 2: the widely used Meta-World benchmark (including ML10 with 10 tasks and ML50 with 50 tasks) [Yu et al., 2020b], 2DMaze, 3DPickup [Chevalier-Boisvert et al., 2024], and MujocoCar [Ji et al., 2023].
+All tasks, including those in Meta-World, are crafted to provide sparse environmental rewards, where the agent receives a reward of $1$ only upon successful completion of the final objective, and 0 otherwise.
 The detailed task configurations are provided in Appendix A.
@@
 (b) 2DMaze: 2D maze tasks where the agent must pick up a key and then pass through a door to exit.
-(b) 3DPickup: 3D maze tasks where the agent aims to navigate to and pick up different target objects at different locations.
-(c) MujocoCar: mujoco-based race car aims to navigate to different specified areas.
+(c) 3DPickup: 3D maze tasks where the agent aims to navigate to and pick up different target objects at different locations.
+(d) MujocoCar: mujoco-based race car aims to navigate to different specified areas.
 5.1 Comparative Evaluation in MTRL
-We benchmark CenRA against several state-of-the-art baselines: (a) the backbone RL algorithms of the policy agents: DQN (Mnih et al.,, 2015) for discrete control tasks and SAC (Haarnoja et al., 2018b, ) for continuous control tasks; (b) the ReLara algorithm (Ma et al., 2024a, ), which can be regarded as a decentralized variant of CenRA, where each policy agent is paired with a separate reward agent, without cross-task information sharing; (c) the Policy Optimization and Policy Correction (PiCor) algorithm (Bai et al.,, 2023); (d) the Multi-Critic Actor Learning (MCAL) algorithm (Mysore et al.,, 2022); (e) the Parameter-compositional MTRL (PaCo) algorithm (Sun et al.,, 2022); and (f) the MTRL with Soft Modularization (SoftModule) (Yang et al.,, 2020).
-They are implemented by either the CleanRL library (Huang et al.,, 2022) or official codebases.
+We benchmark CenRA against several state-of-the-art baselines: (a) the backbone RL algorithms of the policy agents: DQN [Mnih et al., 2015] for discrete control tasks and SAC [Haarnoja et al., 2018b] for continuous control tasks; (b) the ReLara algorithm [Ma et al., 2024a], which can be regarded as a decentralized variant of CenRA, where each policy agent is paired with a separate reward agent, without cross-task information sharing; (c) the TD-MPC2 algorithm [Hansen et al., 2024]; (d) the Contrastive Modules with Temporal Attention (CMTA) algorithm [Lan et al., 2023]; (e) the Policy Optimization and Policy Correction (PiCor) algorithm [Bai et al., 2023]; (f) the Multi-Critic Actor Learning (MCAL) algorithm [Mysore et al., 2022]; (g) the Parameter-compositional MTRL (PaCo) algorithm [Sun et al., 2022]; (h) the Shared-Critic (SC) algorithm [Zhang et al., 2021]; and (i) the MTRL with Soft Modularization (SoftModule) [Yang et al., 2020].
+They are implemented by either the CleanRL library [Huang et al., 2022] or official codebases.
 Each task is trained with 10 different random seeds, and the average results are reported.
@@
 ReLara 0.674 $\pm$ 0.105 0.541 $\pm$ 0.057 0.803 $\pm$ 0.065 0.565 $\pm$ 0.088 429.800 $\pm$ 0.655
+TD-MPC2 0.823 $\pm$ 0.091 0.608 $\pm$ 0.032 0.884 $\pm$ 0.046 0.712 $\pm$ 0.051 505.341 $\pm$ 0.712
+CMTA 0.787 $\pm$ 0.076 0.603 $\pm$ 0.026 0.753 $\pm$ 0.037 0.695 $\pm$ 0.043 480.187 $\pm$ 0.623
 PiCor 0.865 $\pm$ 0.230 0.672 $\pm$ 0.123 0.818 $\pm$ 0.053 0.438 $\pm$ 0.085 437.550 $\pm$ 0.663
@@
 PaCo 0.854 $\pm$ 0.045 0.582 $\pm$ 0.022 0.834 $\pm$ 0.057 0.557 $\pm$ 0.072 421.210 $\pm$ 0.635
+SC 0.556 $\pm$ 0.063 0.354 $\pm$ 0.023 0.798 $\pm$ 0.052 0.687 $\pm$ 0.038 400.254 $\pm$ 0.518
 SoftModule 0.630 $\pm$ 0.042 0.423 $\pm$ 0.057 0.822 $\pm$ 0.076 0.486 $\pm$ 0.055 355.125 $\pm$ 0.594
@@
 By extracting and distributing insights from one task to another, the CRA improves the learning efficiency of individual tasks, highlighting the advantages of integrated knowledge management.
-5.2 What Has the Centralized Reward Agent Learned?
-In this section, we visualize the learned knowledge rewards by the centralized reward agent $\mathcal{A}^{rwd}$ in the 2DMaze environment.
-After training on the four tasks in Section 5.1 of the paper, we let the CRA generate the knowledge rewards for each action in every state and visualize the action direction that yields the maximum rewards, $a^{*}=\operatorname*{arg\,max}_{a}{{\pi^{rwd}}^{*}(s_{i},a)},s_{i}\sim S$, in Figure 5.
-(a) Maze #1
-(b) Maze #2
-(c) Maze #3
-(d) Maze #4
-Figure 5: Visualization of the action yielding the maximum knowledge rewards in the four 2DMaze tasks.
-The shaded areas in the figures represent regions within the real task that the agent cannot reach, as it cannot access the space behind the door without picking up the key.
-However, we forced the agent into these areas for evaluation.
-Outside the shaded regions, we observe that the CRA successfully learned meaningful knowledge rewards.
-Before picking up the key, the agent received the highest reward in the corresponding state when moving towards the key.
-Similarly, after picking up the key, the agent received the highest reward when moving towards the door and the final target.
-This demonstrates that in scenarios where the original environmental rewards are sparse, these detailed knowledge rewards can effectively guide the agent to converge more quickly.
 5.2 Knowledge Transfer to New Tasks
@@
 In ReLara, the reward agent is trained anew without pre-learned knowledge.
-The results are presented in Figure 6 and Table 2.
+The results are presented in Figure 4(a) and Table 2.
 We observe that CenRA with further learning achieved rapid convergence, mainly due to the CRA’s ability to retain previously acquired knowledge while continuing to adapt to new tasks through ongoing optimization.
@@
 However, the CRA provides meaningful dense rewards that significantly accelerate the learning process, even during the initial phases.
-Figure 6: Comparison of the learning performance of CenRA with the baselines in new tasks.
+(a) Comparison of the learning performance of CenRA with the baselines in new tasks in the 2DMaze, 3DPickup and MujocoCar domains.
+(b) Actions yielding the highest knowledge rewards in a new 2DMaze task.
+Figure 4: Experimental results for knowledge transfer to new tasks.
 Table 2: Episodic returns (mean $\pm$ standard error) of all trained agents in the new tasks, tested over 100 episodes in each domain ($\uparrow$ higher is better).
@@
 DQN/SAC 0.228 $\pm$ 0.105 0.210 $\pm$ 0.034 0.263 $\pm$ 0.084 0.158 $\pm$ 0.003 129.055 $\pm$ 0.296
-(a) Before getting the key.
-(b) After getting the key.
-Figure 7: Actions yielding the highest knowledge rewards at each state in a new 2DMaze task.
 To further demonstrate CenRA’s transferability, we select the 2DMaze environment to visualize the knowledge provided by CRA when facing an unseen task.
-As shown in Figure 7, we plot the directions of actions that yield the maximum knowledge reward at each position, categorized into two scenarios: before and after obtaining the key.
+As shown in Figure 4(b), we plot the directions of actions that yield the maximum knowledge reward at each position, categorized into two scenarios: before and after obtaining the key.
 While some guidance in peripheral regions may appear slightly misaligned, most states receive reasonable rewards that align with human understanding.
-This demonstrates the effectiveness of knowledge transfer, and with such dense rewards, the agent’s adaptation to new tasks is able to be well supported.
+This demonstrates the effectiveness of knowledge transfer and with such dense rewards, the agent’s adaptation to new tasks is able to be well supported.
+In addition, to better understand the knowledge learned by the CRA, we provide a case study in Appendix C, where we visualize the CRA-provided rewards.
+This further verifies that the CRA is capable of capturing domain-relevant, task-specific, and semantically meaningful signals across different tasks.
+Table 3: Comparison of CenRA with ablation of different batch sampling control weights.
+Algo. 2DMaze Var. $\downarrow$ ($\times 10^{-2}$)
+Maze #1 Maze #2 Maze #3 Maze #4
+CenRA ($\alpha=0.5$) 0.893 $\pm$ 0.033 0.908 $\pm$ 0.022 0.924 $\pm$ 0.020 0.932 $\pm$ 0.020 0.021
+CenRA ($\alpha=0.25$) 0.889 $\pm$ 0.031 0.901 $\pm$ 0.025 0.915 $\pm$ 0.023 0.925 $\pm$ 0.024 0.065
+CenRA ($\alpha=0.75$) 0.891 $\pm$ 0.030 0.905 $\pm$ 0.024 0.918 $\pm$ 0.021 0.928 $\pm$ 0.022 0.049
+w/o $\bm{w}^{sim}$ ($\alpha=0$) 0.884 $\pm$ 0.033 0.922 $\pm$ 0.021 0.873 $\pm$ 0.041 0.820 $\pm$ 0.039 0.172
+w/o $\bm{w}^{per}$ ($\alpha=1$) 0.758 $\pm$ 0.062 0.884 $\pm$ 0.030 0.824 $\pm$ 0.052 0.867 $\pm$ 0.020 0.284
+w/o both 0.632 $\pm$ 0.053 0.833 $\pm$ 0.041 0.629 $\pm$ 0.08 0.802 $\pm$ 0.054 1.235
+Algo. 3DPickup Var. $\downarrow$ ($\times 10^{-2}$)
+Ball Cube Key Health kit
+CenRA ($\alpha=0.5$) 0.951 $\pm$ 0.020 0.683 $\pm$ 0.090 0.795 $\pm$ 0.062 0.688 $\pm$ 0.067 1.570
+CenRA ($\alpha=0.25$) 0.942 $\pm$ 0.023 0.671 $\pm$ 0.091 0.782 $\pm$ 0.066 0.715 $\pm$ 0.061 1.650
+CenRA ($\alpha=0.75$) 0.938 $\pm$ 0.025 0.665 $\pm$ 0.095 0.775 $\pm$ 0.069 0.702 $\pm$ 0.065 1.723
+w/o $\bm{w}^{sim}$ ($\alpha=0$) 0.822 $\pm$ 0.065 0.702 $\pm$ 0.093 0.704 $\pm$ 0.072 0.887 $\pm$ 0.038 0.892
+w/o $\bm{w}^{per}$ ($\alpha=1$) 0.779 $\pm$ 0.072 0.404 $\pm$ 0.093 0.631 $\pm$ 0.080 0.438 $\pm$ 0.102 3.051
+w/o both 0.811 $\pm$ 0.073 0.457 $\pm$ 0.058 0.483 $\pm$ 0.079 0.370 $\pm$ 0.079 3.796
+Algo. MujocoCar Var. $\downarrow$ ($\times 10^{3}$)
+Target #1 Target #2 Target #3 Target #4
+CenRA ($\alpha=0.5$) 588.221 $\pm$ 0.732 549.337 $\pm$ 0.640 447.743 $\pm$ 0.672 474.320 $\pm$ 0.657 4.244
+CenRA ($\alpha=0.25$) 575.153 $\pm$ 0.740 538.912 $\pm$ 0.655 439.850 $\pm$ 0.680 462.116 $\pm$ 0.665 4.871
+CenRA ($\alpha=0.75$) 580.431 $\pm$ 0.735 542.765 $\pm$ 0.648 441.033 $\pm$ 0.675 468.529 $\pm$ 0.660 4.533
+w/o $\bm{w}^{sim}$ ($\alpha=0$) 319.926 $\pm$ 0.590 486.767 $\pm$ 0.712 332.506 $\pm$ 0.695 260.921 $\pm$ 0.544 9.288
+w/o $\bm{w}^{per}$ ($\alpha=1$) 320.887 $\pm$ 0.891 355.325 $\pm$ 0.677 344.145 $\pm$ 0.872 308.215 $\pm$ 0.723 0.457
+w/o both 57.532 $\pm$ 0.352 257.010 $\pm$ 0.677 677.285 $\pm$ 0.540 77.635 $\pm$ 0.255 82.720
 5.3 Effect of Sampling Weight
-We conducted experiments to understand the effects of the information synchronization mechanism in the CenRA.
-Specifically, we compare the full CenRA model against three variants: (a) CenRA without the similarity weight $\bm{w}^{sim}$, (b) CenRA without the performance weight $\bm{w}^{per}$, and (c) CenRA without the entire sampling weight.
+We conduct experiments to understand the effects of the information synchronization mechanism in the CenRA.
+Specifically, we compare the full CenRA model against five variants: (a) and (b) CenRA with different values of the balance factor $\alpha$, i.e., $\alpha=0.25$ and $\alpha=0.75$, to examine the impact of different weight combinations; (c) CenRA without the similarity weight $\bm{w}^{sim}$ (i.e., $\alpha=0$); (d) CenRA without the performance weight $\bm{w}^{per}$ (i.e., $\alpha=1$); and (e) CenRA without the entire sampling weight.
 To better illustrate the differences among tasks and highlight the role of sampling weights in task coordination and synchronization, we select the four-task domains, i.e., 2DMaze, 3DPickup, and MujocoCar.
-The comparison results are shown in Table 3.
-Table 3: Comparison of CenRA with ablation of different batch sampling control weights.
-Algo. 2DMaze Var. $\downarrow$ (×$\times 10^{-2}$)
-Maze #1 Maze #2 Maze #3 Maze #4
-CenRA 0.89 $\pm$ 0.03 0.91 $\pm$ 0.02 0.92 $\pm$ 0.02 0.93 $\pm$ 0.02 0.02
-w/o $\bm{w}^{sim}$ 0.88 $\pm$ 0.03 0.92 $\pm$ 0.02 0.87 $\pm$ 0.04 0.82 $\pm$ 0.04 0.17
-w/o $\bm{w}^{per}$ 0.76 $\pm$ 0.06 0.88 $\pm$ 0.03 0.82 $\pm$ 0.05 0.87 $\pm$ 0.02 0.28
-w/o both 0.63 $\pm$ 0.05 0.83 $\pm$ 0.04 0.63 $\pm$ 0.08 0.80 $\pm$ 0.05 1.24
-Algo. 3DPickup Var. $\downarrow$ (×$\times 10^{-2}$)
-Ball Cube Key Health kit
-CenRA 0.95 $\pm$ 0.02 0.68 $\pm$ 0.09 0.80 $\pm$ 0.06 0.69 $\pm$ 0.07 1.57
-w/o $\bm{w}^{sim}$ 0.82 $\pm$ 0.07 0.70 $\pm$ 0.09 0.70 $\pm$ 0.07 0.89 $\pm$ 0.04 0.89
-w/o $\bm{w}^{per}$ 0.78 $\pm$ 0.07 0.40 $\pm$ 0.09 0.63 $\pm$ 0.08 0.44 $\pm$ 0.10 3.05
-w/o both 0.81 $\pm$ 0.07 0.46 $\pm$ 0.06 0.48 $\pm$ 0.08 0.37 $\pm$ 0.08 3.80
-Algo. MujocoCar Var. $\downarrow$ (×$\times 10^{3}$)
-Target #1 Target #2 Target #3 Target #4
-CenRA 588.2 $\pm$ 0.73 549.3 $\pm$ 0.64 447.7 $\pm$ 0.67 474.3 $\pm$ 0.66 4.24
-w/o $\bm{w}^{sim}$ 319.9 $\pm$ 0.59 486.8 $\pm$ 0.71 332.5 $\pm$ 0.70 260.9 $\pm$ 0.54 9.29
-w/o $\bm{w}^{per}$ 320.9 $\pm$ 0.89 355.3 $\pm$ 0.68 344.1 $\pm$ 0.87 308.2 $\pm$ 0.72 0.46
-w/o both 57.5$\pm$0.35 257.0 $\pm$ 0.68 677.3 $\pm$ 0.54 77.6 $\pm$ 0.26 82.7
+The results are shown in Table 3.
 The results indicate that the two weights, which control the allocation of samples drawn from each policy agent’s experiences, mainly influence the overall learning performance.
 Specifically, the absence of sampling weight leads to unbalanced learning outcomes, which is observed by the increased variance in episodic returns across four tasks.
+In contrast, when both weights are incorporated, the learning process becomes notably more stable, indicating that the joint consideration of task similarity and learning progress is essential for coordinated optimization.
 While the full CenRA model does not always achieve the lowest variance, it consistently outperforms the other three ablation models regarding overall system performance.
-Both weights play essential roles in the information synchronization, with the performance weight $\bm{w}^{per}$ having a more important impact.
-It allows CRA to pay more attention to the policy agents that are underperforming or progressing slowly.
-This suggests that considering the overall learning performance of the multi-task system is also an important objective that CenRA seeks to achieve.
+Both weights play essential roles in information synchronization, with the performance weight $\bm{w}^{\mathrm{per}}$ having a more significant impact.
+It allows the CRA to focus more on policy agents that are underperforming or progressing slowly, ensuring balanced system-wide learning.
+Moreover, different choices of the balancing factor $\alpha$ emphasize distinct aspects of synchronization: a larger $\alpha$ highlights task similarity and promotes uniformity across related tasks, whereas a smaller $\alpha$ prioritizes compensating lagging tasks by amplifying the effect of $\bm{w}^{\mathrm{per}}$.
+This flexible weighting further enhances stability and adaptability, demonstrating that considering the overall learning performance of the multi-task system is a central objective that CenRA seeks to achieve.
 6 Discussion and Conclusion
-We propose a novel framework named CenRA that innovatively integrates reward shaping into MTRL.
-The framework not only shares domain knowledge across tasks to improve learning efficiency but also effectively addresses the sparse-reward challenge.
-Specifically, the centralized reward agent functions as a knowledge pool, responsible for knowledge distillation and distribution across multiple tasks.
-Furthermore, our information synchronization mechanism mitigates knowledge distribution disparities across tasks, ensuring optimal system-wide performance.
-Experiments demonstrate that the dense knowledge rewards generated by the CRA efficiently guide the policy agents’ learning, leading to faster convergence compared to baseline methods.
-The CenRA also performs superiorly and robustly in transferring acquired knowledge to new tasks.
-CenRA’s current limitation is its requirement for uniform state and action shapes across tasks.
-Future improvements could explore preprocessing techniques to adapt the framework for varying state and action shapes to broaden CenRA’s applicability.
+We propose a novel framework CenRA that integrates reward shaping into multi-task reinforcement learning.
+The framework shares domain knowledge across tasks to improve learning efficiency and effectively addresses the sparse-reward challenge.
+Specifically, the centralized reward agent (CRA) functions as a knowledge pool, responsible for distilling and distributing knowledge across tasks.
+Furthermore, the information synchronization mechanism mitigates imbalances in knowledge distribution, ensuring optimal system-wide performance.
+Experiments demonstrate that dense knowledge rewards generated by the CRA effectively guide policy learning, leading to faster convergence than baseline methods.
+CenRA also demonstrates superior and robust transferability to new tasks.
+CenRA’s main limitation is its requirement for consistent state and action dimensions across tasks.
+Future work could explore preprocessing techniques to adapt the framework to varying task structures, broadening its applicability.
 Additionally, the fixed trade-off between similarity weight and performance weight may not be ideal.
-A more flexible approach, potentially through adaptive regulation of these weights, could further enhance the framework.
-Besides, the performance weight might favor underperforming tasks, achieving overall balance, but potentially limiting the upper bound of top-performing tasks.
-This issue could necessitate a better mechanism for trade-offs.
+A more flexible approach, such as adaptive weight regulation, could further enhance the framework.
+Moreover, the performance weight might favor underperforming tasks to achieve overall balance, but could limit the performance ceiling of high-performing tasks, indicating the need for a more effective trade-off mechanism.
+Acknowledgments and Disclosure of Funding
+This work was supported by an Academic Research Grant MOE-T1 251RES2408 and a Research Scholarship from the Ministry of Education, Singapore.
 Appendix A Mutli-Task Experimental Configurations
@@
 The detailed configurations of each task are illustrated in Figure 5.
-The Meta-World tasks illustration is adapted from (Yu et al., 2020b, ).
+The Meta-World tasks illustration is adapted from [Yu et al., 2020b].
 Figure 5: Illustration of multiple tasks in different domains in our experiments.
-Appendix B Experimental Configurations
+Appendix B Network Structures and Hyperparameters
 B.1 Network Structures
@@
 Figure 6: The structures of neural networks in our implementation.
-B.2 Computing Resources
+B.2 Hyperparameters
+We have observed that CenRA demonstrated high robustness and was not sensitive to hyperparameter choices. Table 4 shows the hyperparameters we used in all the experiments.
+Table 4: The hyperparameters of CenRA used in our experiments.
+Module Hyperparameters Values
+Centralized Reward Agent $\mathcal{A}^{rwd}$ discounted factor $\gamma$ 0.99
+batch size 256
+actor module learning rate $3\times 10^{-4}$
+critic module learning rate $1\times 10^{-3}$
+policy networks update frequency (steps) 22
+target networks update frequency (steps) 11
+target networks soft update weight $\tau$ $5\times 10^{-3}$
+burn-in steps 50005000
+Policy Agent $\mathcal{A}^{pol}_{i}$ (DQN Agent) knowledge reward weight $\lambda$ 0.5
+discounted factor $\gamma$ 0.99
+replay buffer size $|\mathcal{D}_{i}|$ $1\times 10^{6}$
+batch size 128
+burn-in steps 1000010000
+Policy Agent $\mathcal{A}^{pol}_{i}$ (SAC Agent) knowledge reward weight $\lambda$ 0.5
+discounted factor $\gamma$ 0.99
+replay buffer size $|\mathcal{D}_{i}|$ $1\times 10^{6}$
+batch size 256
+actor module learning rate $3\times 10^{-4}$
+critic module learning rate $1\times 10^{-3}$
+SAC entropy term factor $\alpha$ learning rate $1\times 10^{-4}$
+policy networks update frequency (steps) 22
+target networks update frequency (steps) 11
+target networks soft update weight $\tau$ $5\times 10^{-3}$
+burn-in steps 1000010000
+B.3 Computing Resources
 The experiments in this paper were conducted on a computing cluster, with the detailed hardware configurations listed in Table 5.
@@
 Brand Supermicro 2022
-B.3 Hyperparameters
-We have observed that CenRA demonstrated high robustness and was not sensitive to hyperparameter choices. Table 5 shows the hyperparameters we used in all the experiments.
-Table 5: The hyperparameters of CenRA used in our experiments.
-Module Hyperparameters Values
-Centralized Reward Agent $\mathcal{A}^{rwd}$ discounted factor $\gamma$ 0.99
-batch size 256
-actor module learning rate $3\times 10^{-4}$
-critic module learning rate $1\times 10^{-3}$
-policy networks update frequency (steps) $2$
-target networks update frequency (steps) $1$
-target networks soft update weight $\tau$ $5\times 10^{-3}$
-burn-in steps $5000$
-Policy Agent $\mathcal{A}^{pol}_{i}$ (DQN Agent) knowledge reward weight $\lambda$ 0.5
-discounted factor $\gamma$ 0.99
-replay buffer size $|\mathcal{D}_{i}|$ $1\times 10^{6}$
-batch size 128
-burn-in steps $10000$
-Policy Agent $\mathcal{A}^{pol}_{i}$ (SAC Agent) knowledge reward weight $\lambda$ 0.5
-discounted factor $\gamma$ 0.99
-replay buffer size $|\mathcal{D}_{i}|$ $1\times 10^{6}$
-batch size 256
-actor module learning rate $3\times 10^{-4}$
-critic module learning rate $1\times 10^{-3}$
-SAC entropy term factor $\alpha$ learning rate $1\times 10^{-4}$
-policy networks update frequency (steps) $2$
-target networks update frequency (steps) $1$
-target networks soft update weight $\tau$ $5\times 10^{-3}$
-burn-in steps $10000$
+Appendix C What Has the Centralized Reward Agent Learned?
+In this section, we visualize the learned knowledge rewards by the centralized reward agent $\mathcal{A}^{rwd}$ in the 2DMaze environment.
+After training on the four tasks in Section 5.1 of the paper, we let the CRA generate the knowledge rewards for each action in every state and visualize the action direction that yields the maximum rewards, $a^{*}=\operatorname*{arg\,max}_{a}{{\pi^{rwd}}^{*}(s_{i},a)},s_{i}\sim S$, in Figure 7.
+The shaded areas in the figures represent regions within the real task that the agent cannot reach, as it cannot access the space behind the door without picking up the key.
+However, we forced the agent into these areas for evaluation.
+Outside the shaded regions, we observe that the CRA successfully learned meaningful knowledge rewards.
+Before picking up the key, the agent received the highest reward in the corresponding state when moving towards the key.
+Similarly, after picking up the key, the agent received the highest reward when moving towards the door and the final target.
+This demonstrates that in scenarios where the original environmental rewards are sparse, these detailed knowledge rewards can effectively guide the agent to converge more quickly.
+(a) Maze #1
+(b) Maze #2
+(c) Maze #3
+(d) Maze #4
+Figure 7: The actions yielding the maximum knowledge rewards in the four 2DMaze tasks.